### v2

# Socialized Learning: A Survey of the Paradigm Shift for Edge Intelligence in Networked Systems

Abstract
Amidst the robust impetus from artificial intelligence (AI) and big data, edge intelligence (EI) has emerged as a nascent computing paradigm, synthesizing AI with edge computing (EC) to become an exemplary solution for unleashing the full potential of AI services. Nonetheless, challenges in communication costs, resource allocation, privacy, and security continue to constrain its proficiency in supporting services with diverse requirements. In response to these issues, this paper introduces socialized learning (SL) as a promising solution, further propelling the advancement of EI. SL is a learning paradigm predicated on social principles and behaviors, aimed at amplifying the collaborative capacity and collective intelligence of agents within the EI system. SL not only enhances the system’s adaptability but also optimizes communication, and networking processes, essential for distributed intelligence across diverse devices and platforms. Therefore, a combination of SL and EI may greatly facilitate the development of collaborative intelligence in the future network. This paper presents the findings of a literature review on the integration of EI and SL, summarizing the latest achievements in existing research on EI and SL. Subsequently, we delve comprehensively into the limitations of EI and how it could benefit from SL. Special emphasis is placed on the communication challenges and networking strategies and other aspects within these systems, underlining the role of optimized network solutions in improving system efficiency. Based on these discussions, we elaborate in detail on three integrated components: socialized architecture, socialized training, and socialized inference, analyzing their strengths and weaknesses. Finally, we identify some possible future applications of combining SL and EI, discuss open problems and suggest some future research, with the hope of arousing the research community’s interest in this emerging and exciting interdisciplinary field.
Index Terms:
Edge intelligence (EI), socialized learning (SL), artificial intelligence (AI), edge computing (EC)
I Introduction
With the rapid advancement of information technology and the proliferation of wireless communications and Internet of Things (IoT) devices, human society is confronted with both the challenges and opportunities presented by vast quantities of data [1]. Data serves not only as a critical resource but also as a potent catalyst, propelling progress and innovation in the field of artificial intelligence (AI) [2]. AI has emerged as a transformative technology in today’s world, finding extensive application across diverse domains, including healthcare, education, transportation, manufacturing, and entertainment, among others [3]. However, AI’s evolution imposes heightened demands on data processing, necessitating substantial bandwidth and robust computational resources to achieve efficient, precise, and real-time data analysis and decision-making.
I-A From Cloud Computing to Edge Intelligence (EI)
Traditional cloud computing sends data from devices like sensors and smartphones to the cloud for processing, then returns results to users. Although this approach can provide abundant computational and storage resources, it also includes difficulties such as latency in the network, bandwidth usage, and privacy and security issues. In response to these issues, a novel computing paradigm, edge computing (EC), has acquired extensive attention in recent years [4]. EC shifts computational services to the network edge, reducing data transmission and enhancing processing efficiency and user experience. The transition from centralized cloud computing to EC underscores the necessity of robust communication frameworks, network design, and resource allocation, catering to the unique demands of edge environments for efficient system operation and data transmission.
Figure 1: Social learning in human society versus SL in EC.
Building upon the foundation of EC, EI has been introduced as an emergent paradigm, integrating AI with EC to facilitate intelligent data processing and decision-making at the network edge [5, 6, 7]. EI, employing AI algorithms and models on edge devices, can rapidly perceive and train on vast troves of local data [8]. This enables rapid adaptation to evolving environments, reduced latency, and enhanced computing efficiency. Consequently, EI is regarded as a formidable driving force in the future development of networks. EI’s reliance on AI algorithms at the network edge also amplifies the need for efficient communication, computing, resource allocation, and other aspects to manage complex data flows and interactions among various devices.
Primarily, EI can fulfill critical requirements such as agile connectivity, real-time operations, data optimization, intelligent applications, and security and privacy [9]. For instance, in autonomous driving scenarios, EI enables rapid communication and collaborative control between vehicles or between vehicles and roadside infrastructure, offering timely responses in emergencies [10]. The collaboration of communication, networking, computing, resource, and system optimization plays a pivotal role in ensuring that EI operates at peak efficiency. Furthermore, EI is anticipated to reach a phase of stability in the coming years and find widespread application in numerous burgeoning domains. For example, in the metaverse, EI can utilize artificial intelligence-generated content (AIGC) technologies on edge devices to construct and render virtual worlds, providing immersive experiences in virtual reality (VR) and augmented reality (AR).
I-B Combination of EI and Socialized Learning (SL)
Despite the immense potential and prospects of EI, its evolution still confronts numerous challenges and dilemmas [11]. For instance, users inevitably generate analogous machine learning (ML) tasks, potentially necessitating identical categories of data or even anticipating congruent training outcomes. This massive repetition of model training can lead to a substantial redundancy in computations across the network and a squandering of finite edge resources. Moreover, smaller data sizes can lead to model overfitting. Consequently, prevalent meaningless model training results in diminished training efficiency and a degradation in quality of service (QoS).
Another challenge encompasses achieving efficient communication, resource allocation, and privacy protection within EI. Given the involvement of various types and devices, pivotal issues that demand resolution include coordinating collaboration and competition among them, balancing mutual interests and costs, and ensuring security in interactions [12].
To address these challenges, the concept of SL is introduced as an innovative solution. As shown in Fig. 1, SL is the adaptation or mapping of social learning into machine interactions [13]. It is aimed at creating a cohesive machine society where social rules are implemented to facilitate collaborative interactions among devices, thereby enhancing overall system intelligence. SL is inspired by the concept of social learning in human societies. In these societies, individuals acquire knowledge, skills, attitudes, and values through observation, imitation, communication, and feedback, integrating these learnings into their behavior and thought processes.
Figure 2: The taxonomy graph on the integration of EI and SL.
Hence, SL is a learning paradigm based on social principles and behaviors, aimed at enhancing the collaborative capabilities and collective intelligence of various agents within the EI system. SL draws inspiration from a myriad of social interaction modalities present in human societies, such as cooperation, submissive and conflict interactions, and applies them within the EI system. SL perceives each device as an independent, autonomous entity with social awareness. These devices can learn not only from their own data and experiences but also from the data and experiences of others. Through this method of social interaction, SL can effectively utilize data resources within the EI, reduce redundant computations, and enhance training efficiency and QoS. Moreover, SL can effectively allocate computational resources within the EI system, optimize communication costs, and safeguard privacy and security. SL promises to endow the EI system with greater intelligence, adaptability, and robustness.
A real-life application of SL in EI. In the context of intelligent transportation systems, the integration of EI and SL enables autonomous vehicles to learn collaboratively and adapt to dynamic road conditions. A network of self-driving vehicles, equipped with EI capabilities and SL algorithms, continuously gathers and processes real-time data about traffic, road conditions, and weather using onboard sensors and EC resources. This decentralized approach allows for swift responses to immediate situations [14]. The true potential of this system lies in the application of SL principles, enabling vehicles to learn from their own experiences and the collective wisdom of the entire system. When a vehicle encounters a novel situation, such as an unexpected obstacle or congestion, it shares this information with other vehicles in the network [15]. This shared knowledge is incorporated into the collective intelligence, benefiting all vehicles.
For instance, if a vehicle detects a hazard, it can alert other vehicles to take alternate routes or proceed with caution. This real-time information sharing contributes to the long-term learning and optimization of the entire transportation network. As more vehicles encounter and share similar situations, the system develops a comprehensive understanding of the city’s traffic dynamics and potential hazards [16]. This collaborative learning approach allows vehicles to proactively adapt to changing conditions based on insights from the fleet. Over time, this learning process enhances the safety, efficiency, and resilience of the autonomous transportation system, leading to smoother traffic flow, reduced congestion, and minimized accidents. The combination of EI and SL exemplifies the transformative potential of these technologies in revolutionizing transportation networks. By leveraging EC and SL, vehicles make intelligent decisions in real time while contributing to the collective intelligence of the system, ultimately improving urban transportation.
I-C Scenarios Overview of SL in EI
The complex synergy of EI and SL has been integrated and reflected in various fields. This fusion leverages the strengths of both EI and SL, enabling systems to process data at the edge efficiently and enhance system intelligence. Additionally, it allows systems to learn and adapt through shared experiences and social principles, strengthening their adaptability to complex environments and user experiences.
Across diverse application scenarios such as autonomous driving, IoT, smart healthcare, recommendation systems, smart transportation, and smart cities, the integration of SL in EI exhibits certain common characteristics. These include collaborative learning among edge devices, improved adaptability to dynamic conditions, and enhanced privacy and security due to the decentralized nature of SL. However, the specific objectives and implementation details of SL in EI may vary depending on the unique requirements and constraints of each application domain. In the following contents, we will briefly explore how the integration of SL and EI manifests in various current application scenarios, showcasing the potential of this powerful combination to drive innovation and transform diverse sectors. A more detailed discussion on the future applications of SL in EI and a comparative analysis of these applications will be provided in Section VIII-A.
Autonomous Driving. The integration of EI and SL primarily manifests in the use of edge devices for real-time data processing and decision-making. The deployment on vehicles enables collaborative learning among vehicles, sharing road condition information, thereby enhancing overall driving safety and efficiency. Vehicles can react based on their own sensory inputs and learn from other vehicles’ experiences through communication and data sharing.
IoT. The integration of EI and SL enables devices to process data and share knowledge and experiences through SL algorithms. In smart home systems, this facilitates accurate environmental control and user behavior prediction, elevating IoT system intelligence through device interconnectivity and data sharing.
Smart Healthcare. The combination of EI and SL improves efficiency and accuracy in medical data processing. By deploying edge resources for preliminary data processing and analysis, the burden on central servers is alleviated, while patient data privacy is protected. Edge-assisted medical systems allow for cost savings and improved healthcare service delivery, facilitated by shared knowledge and experiences among medical devices.
Recommendation Systems. EI and SL integration is exemplified in the use of EC for personalized recommendations and user data processing, safeguarding privacy. Personalization is achieved through edge-based user behavior analysis and model training, ensuring privacy while offering tailored recommendations. SL algorithms further optimize the system based on user interactions and feedback.
Smart Transport and Unmanned Aerial Vehicles (UAVs). EI and SL integration is evident in the use of UAVs for data collection and processing. Deploying EC and SL algorithms on UAVs facilitates real-time monitoring and data processing, improving system response time and efficiency. UAVs sharing flight experiences and data processing strategies through SL enhance the performance of the entire system.
Smart City. The convergence of EI and SL aids in more effective city data management and processing. Deploying edge devices across different city areas and utilizing SL for data and experience sharing facilitate intelligent and automated city management. For instance, data sharing and collaborative learning among traffic monitoring devices can lead to more effective traffic flow management and accident prevention.
The above examples provide a concise overview of how SL and EI are currently integrated across various application domains, highlighting the common characteristics and potential benefits of this synergistic combination. However, it is important to note that this section serves as a brief introduction to the current state of SL in EI applications. A more comprehensive analysis of the future applications of SL in EI, along with a detailed comparative study of these application scenarios, will be presented in Section VIII-A.
I-D Our Contributions
This article aims to explore the integration of SL with EI and offers practical tutorials. More specifically, we enumerate the following principal contributions to this paper:
  • We first outline the basic definition and architecture of EI and SL and discuss the necessity of SL in the presence of EI. We also present the problems addressed by EI.
  • We discuss the motivations for combining SL and EI from two perspectives. SL can be instrumental in addressing several challenges faced by EI, thereby compensating for the shortcomings of conventional methods. Further, the integration of SL within EI plays a pivotal role across various facets, optimizing system performance.
  • We summarize three ideas for deploying SL training and inference tasks in the EI architecture based on existing studies and analyze their advantages and disadvantages.
  • We also review the future applications for applying SL to other fields under the EI architecture. Additionally, we identify some open problems and discuss future research on combining SL and EI.
This paper’s structure, depicted in Fig. 2, commences with an introduction to the backgrounds of EI and SL, followed by a review of relevant concepts in Sec. II. The rationale for integrating EI and SL is in Sec. III, followed by an exposition on the integration points: socialized architecture, training, and inference in Secs. IVVI. Tutorials of SL in EI are presented in Sec. VII. Applications, open problems, and future research are discussed in Sec. VIII, concluding with Sec. IX.
II Background and fundamentals of EI and SL
II-A EI: Definition, Challenges, and Solutions
The surge in connected devices and AI capabilities has strained cloud infrastructure for real-time applications. EI, characterized by the integration of computation at the network’s edge, offers a viable solution to these evolving challenges. We summarize the important definitions and related abbreviations listed in Table I.
II-A1 Definition of EI
EI is the convergence of AI and EC, as detailed below:
  • AI: AI simulates human intelligence, encompassing techniques like deep learning (DL). The main objective of AI is to model intelligent human behavior by discerning patterns and insights from extensive data repositories.
  • EC: This refers to a computational paradigm that extends beyond traditional cloud architectures. It offers computational services closer to the data sources or the ‘edge’ of the network [17]. The ‘edge’ encompasses a spectrum of resources ranging from data origins to cloud-based centers, predominantly excluding the cloud itself. Unlike fog computing, which has an inherent hierarchical structure, EC emphasizes the significance of individual nodes that may not necessarily be interconnected in a network [18].
  • Interplay of AI and EC: EI emerges naturally from the synergy of AI and EC. AI provides advanced techniques, while EC supplies varied, real-time application platforms [8]. EC’s processing at the data source reduces latency, which is critical for analyzing large datasets such as audio and video. These rapid insights power applications like public transport planning [19], urban monitoring [20] and early forest fire alerts.
TABLE I: IMPORTANT ABBREVIATIONS LIST IN ALPHABETICAL ORDER
Abbr. Definition Abbr. Definition Abbr. Definition
AI artificial intelligence EI edge intelligence IoT Internet of Things
EC edge computing FL federated learning ES edge server
CPU central processing unit LAN local area network QoS quality of service
ShI shared intelligence DI distributed intelligence MAS multi-agent system
CI collective intelligence ML machine learning D2D device-to-device
KD knowledge distillation QoE quality of experience RL reinforcement learning
GAN generative adversarial network DNN deep neural network MEC mobile edge computing
DL deep learning MARL multi-agent reinforcement learning CrI crowd intelligence
DA domain adaptation MTL multi-task learning NLP natural language processing
HRL hierarchical reinforcement learning DAG directed acyclic graph CPN computing power network
AIGC artificial intelligence-generated content AR augmented reality VR virtual reality
SL socialized learning SI swarm intelligence DRL distributed reinforcement learning
MPC multi-party computation SMMA secure multi-party model aggregation IRS intelligent reflecting surface
II-A2 Challenges in Implementing EI
EI offers a new approach to data processing but faces the following challenges:
  • Trustworthiness: Privacy, security, and regulatory compliance, such as general data protection regulation (GDPR), are crucial in EI’s decentralized framework, which heightens exposure to security risks [21].
  • Scalability: The limited computational and storage capacities of edges can hinder the efficient deployment of AI models. Striking a balance between latency, energy consumption, and model accuracy becomes a critical concern in EI.
  • Resource Constraints: Edges are often constrained in terms of computation, storage, and communication bandwidth. It can be difficult to effectively manage these resources while maintaining the best possible AI performance [22].
II-A3 Solutions to Overcome Challenges
To address the aforementioned challenges, the research community and industry have proposed several innovative solutions:
  • Advanced Learning Models and Architectures: Techniques such as federated learning (FL), swarm intelligence (SI), and split learning are used to enhance data privacy and training efficiency at the edge [23, 24]. Methods such as reinforcement learning (RL) offer dynamic adaptability in adversarial environments [25, 26].
  • Security Enhancements: Incorporating cryptographic techniques like differential privacy and Lagrange-coded computing can strengthen EI architectures against potential breaches. The use of quantum computing and blockchain further ensures data integrity and confidentiality [27].
  • Task-oriented Communication: A shift from data-centric communication approaches to task-oriented paradigms can address the scalability concerns. By focusing on optimizing tasks over maximizing data rates, networks can efficiently handle the demands of EI applications.
The role of advanced communication protocols, robust networking strategies, efficient computing, and resource allocation becomes increasingly pivotal, especially for ensuring efficient collaboration. By understanding its challenges and actively developing innovative solutions, EI can reshape the landscape of communications and AI deployments.
II-B The Basic Operation Rules in Society
There are three basic forms of social interaction: cooperative, confrontational, and submissive interactions [28]. These interactions are crucial in shaping collective evolution. Each form contributes uniquely to the complex structure of interpersonal relationships, reflecting the intricacies and diverse outcomes of our social environment.
II-B1 Submissive Interaction
Submissive interactions, characterized by one party yielding to another, play a crucial role in achieving social equilibrium, evident across diverse settings like workplaces, families, and educational institutions. This dynamic, while promoting social stability [28], may also inadvertently lead to power imbalances. Suggestion is a subtle but pervasive form of influence where individuals or entities use authority or trust to shape the actions or beliefs of others. Imitation serves as a critical mechanism for social relations, reflecting observed behaviors either deliberately or innately, thereby bridging the gap between external actions and personal behaviors [29]. This process can enhance social cohesion but also risks reinforcing undesirable actions. Meanwhile, conformity emerges from the pursuit of social acceptance [30], manifesting in behaviors ranging from strict adherence to social rules to more subtle adjustments in conduct.
II-B2 Cooperation Interaction
Cooperative interaction is fundamental in shaping an interconnected world, representing the amalgamation of individual or group efforts towards common goals, thereby fostering social progress. At its core, cooperation involves voluntary coordination of resources, knowledge, and skills, emphasizing synergistic collaboration for shared objectives or mutual interests without direct resource exchange [31]. In contrast, exchange is defined by the principle of reciprocity, focusing on a balanced give-and-take of resources, information, or services to preserve equitable relationships and trust [32]. This differs from assistance, which is characterized by one-way support provided without the expectation of immediate return, often driven by altruism, empathy, or a sense of social responsibility [33]. Lastly, mutual assistance combines elements of cooperation, exchange, and assistance, highlighting bidirectional support and shared responsibility based on interdependence among participants [34].
II-B3 Conflict Interaction
Conflict interaction, in contrast to cooperative dynamics, encompasses disagreements and rivalries arising from divergent beliefs or interests, ranging from subtle disagreements to open conflicts. Such interactions can be a double-edged sword. Competition, as a form of conflict interaction, involves entities competing for limited resources or recognition. While it can drive innovation and excellence, unbalanced competition may lead to discord and inequity [35]. Compulsory interaction, on the other hand, is characterized by one party imposing its will on another, creating a power dynamic that can maintain social order or, if unchecked, may lead to resentment and power imbalances [36]. The third subtype, struggle, represents a more intense form of conflict where entities actively contend over differing beliefs, values, or interests. Although initially disruptive, struggles can signify underlying needs for change, potentially leading to long-term social improvements [37].
II-B4 Deciphering SL-emergence and Essence
At its core, SL borrows heavily from the rules that govern human social interaction, rendering it akin to an intelligent society.
SI: A Preliminary Analogy. SI arises from the collective behavior of decentralized, self-organized systems often inspired by natural phenomena, like ant colonies or bird flocks [38]. It is characterized by local interactions among individuals without centralized control. For example, in an ant colony, tasks such as nest building, larvae care, defense, and foraging are distributed without centralized command, despite a queen’s presence. Ants rely on local communication, like pheromone trails, for complex problem-solving and colony survival [39].
From Swarm to Society: The Leap to SL. While SI is rooted in mimicking natural systems, SL moves a step further by integrating principles inherent in human society. SL is concerned with both artificial and natural systems wherein multiple entities coordinate using self-organization and decentralized control, much like SI. However, SL distinguishes itself by more heavily drawing from the norms, behaviors, and interactions prevalent in human society.
SL: Social Norms Meeting Intelligence. At its core, SL is about emergent collective intelligence. It encapsulates systems that encompass multiple independent devices whose behaviors and interactions are reminiscent of human social operations. By integrating the lessons from human society, SL aims to address real-world challenges in more sophisticated, nuanced ways, transcending the capabilities of traditional SI systems.
While SI offers an illustrative backdrop to understand the decentralized and self-organizing aspects of SL, it is the addition of human social principles that truly defines and differentiates SL. This integration, drawing from the richness of social norms and behaviors, allows SL to offer more holistic and adaptive solutions. Furthermore, the implementation of SL in EI systems necessitates sophisticated computing techniques and system designs to facilitate effective communication and coordination among devices, optimizing system performance.
The relationship between SL and SI. Although SL and SI share fundamental principles like decentralized control and self-organization, they differ in scope and complexity. SI focuses on simple, local interactions among homogeneous agents, while SL incorporates diverse social norms, behaviors, and heterogeneous interactions found in human societies. SL emphasizes individual autonomy and social awareness in shaping collective behavior, enabling it to tackle complex, real-world challenges in EI systems. SI, in contrast, may be more applicable to specific EI scenarios involving homogeneous devices or simple coordination tasks. By leveraging social principles and learning mechanisms, SL fosters the development of socially-aware, adaptive, and collaborative EC systems, addressing a broader range of challenges compared to traditional SI approaches.
II-B5 Formal Definition of SL
SL is a paradigm that integrates the principles and dynamics of human social interactions into the learning processes of intelligent systems. It aims to create a cohesive and adaptive learning environment where multiple agents, such as devices or algorithms, can learn collaboratively and share knowledge, experiences, and resources, guided by social norms and behaviors. SL draws inspiration from various social interaction modalities, including cooperation, competition, and hierarchical relationships, to optimize learning outcomes, resource allocation, and collective decision-making. By incorporating social principles, SL enables systems to tackle complex, real-world challenges more effectively, adapting to dynamic environments and leveraging the collective intelligence of the participating agents.
II-C SL: A Function incorporating Multiple Intelligence
Figure 3: The timelines of related concepts.
TABLE II: Comparison of Various Types of Concepts
Concept
Entity
Type
Representative
Algorithms
Decision
Structure
Main
Advantages
Main
Challenges
Result
Sharing
Wisdom of
the crowd [40]
People, esp.,
non-experts
Aggregation methods,
Diversity-based methods
Distributed
Collective judgment,
error cancellation
Independent opinions,
unbiased judgment
$\checkmark$
Shared
Intelligence [41]
Humans,
Computers
Collaborative tools,
Content management Systems
Distributed
Effective knowledge
utilization, enhanced
collaboration
Trust,
privacy issues
$\checkmark$
Distributed
Intelligence [42]
Humans,
Computers
Distributed computing,
distributed gradient descent
Distributed
Scalability,
fault tolerance
Coordination,
communication
overhead
$\checkmark$
Swarm
Intelligence [38]
Robots,
Algorithms
Particle Swarm Optimization,
Ant Colony Optimization
Distributed
Solves complex problems,
good adaptability
Coordination,
scalability issues
$\times$
Multi-agent
Systems [43]
AI Agents
Coordination Graphs,
Q-Learning
Distributed
Solves complex problems,
good resilience
Coordination,
collaboration issues
$\checkmark$
Collective
Intelligence [44]
Humans,
Computers
Social Network Analysis,
Collaborative Filtering
Distributed
Strong innovation,
rich knowledge
Quality control,
coordination issues
$\checkmark$
ML
Ensembles [45]
ML Models
Random Forests,
Boosting
Ensemble
High accuracy,
good robustness
High computational,
storage requirements
$\checkmark$
Crowd
Intelligence [46]
People, esp.,
online users
Crowdsourcing algorithms,
Reputation systems
Distributed
Solves complex problems,
high diversity
Quality control,
incentive issues
$\checkmark$
We will introduce the function form of SL. The goal is to adjust different variables of function to achieve different SL schemes that can adapt to various application scenarios and requirements in EI. The SL function can be represented as:
$F(\mathbf{A},\mathbf{R}_{w},\mathbf{R}_{a},\mathbf{I},\mathbf{L}),$ (1)
where $\mathbf{A}$ represents the architecture of the network and other related information. $\mathbf{R}_{w}$ and $\mathbf{R}_{a}$ represent the methods of showing relationships within and across layers, respectively. $\mathbf{I}$ is an identifier that distinguishes inference and training operations. $\mathbf{L}$ indicate the specific operation object.
We can choose appropriate variable values according to different application scenarios and requirements to achieve different SL schemes. For example:
  • If we want to implement FL in EI, we can set $\mathbf{A}$ as a two-layer structure, where the first layer is a server and the other is a set of devices. We can set $\mathbf{R}_{a}$ as cooperation, where each device cooperates with others in the same layer to share models. $\mathbf{I}$ sets as a training operation, and $\mathbf{L}$ sets as DNNs.
  • If we want to implement edge-device collaborative inference in EI, we define a two-layer structure with edge servers (ESs) and devices. $\mathbf{R}_{a}$ is set to assistance, signifying a dynamic where one layer offers or seeks help from another. Inference operations are denoted as $\mathbf{I}$, and tasks as $\mathbf{L}$.
With this function form, we can flexibly express and design various SL schemes and match them with EI. A visual summary of these schemes and their corresponding parameter values will be presented in the taxonomy graph in the end of Section VI, as shown in Fig. 37. This flexibility also extends to the optimization of network architectures and computational resource allocation, ensuring that the integration of SL and EI is not only intelligent but also systemically efficient and communication-coherent.
II-D Opportunities Brought by the Integration of EI and SL
Several recent surveys have investigated the integration of EC with advanced technologies. Wang et al. [47] discussed the convergence of EC and DL, including application scenarios and enabling technologies. Taleb et al. [48] examined 5G network edge cloud architecture. Yang et al. [49] explored the integration of blockchain with EC, addressing key motivations and challenges. Lim et al. [15] reviewed FL in EC, focusing on communication and privacy challenges. Lastly, Ning et al. [20] analyzed EC and ML applications for the Internet of UAVs. While these surveys offer valuable insights, they do not organize and present state-of-the-art research in EI from an SL perspective, providing an innovative and comprehensive understanding of the challenges, techniques, and applications in EI systems.
SL deepens EI’s understanding of cooperation and improves EI systems’ intelligence and collaboration through inspiration from human social interaction. This understanding fosters improved decision-making, addressing gaps that traditional AI might overlook. In decentralized EI, SL harmonizes entities to optimize resources for collective goals. Furthermore, SL’s emphasis on consensus and collective intelligence yields adaptable, resilient systems. This boosts trust and ensures efficient configurations in EI systems.
Conversely, EI’s resources and intelligence offer fertile ground for SL. EI’s distributed architecture enables seamless SL integration. SL entities leverage edge resources to optimize computation and resource allocation. SL models in edge contexts access EI capabilities, creating unique, society-inspired services. Lastly, integrating EI with SL enhances compatibility, scalability, and security, boosting performance.
The integration further facilitates a more dynamic resource allocation and network management, enabling the EI system to adapt and respond to varying computational demands and communication needs, thereby exemplifying a comprehensive system-level approach to modern AI and EC challenges. This survey emphasizes the synergistic potential of EI and SL, aiming to construct a potent platform for maximizing the value of information, stimulating resource and intelligence sharing at the edge, and enhancing system efficiency.
II-E Review of Existing Similar Concepts
In the dynamic AI area, SL emerges as a distinctive advancement, integrating aspects of various established paradigms. Illustrated in Fig. 3 and Table II, SL’s evolution builds upon these concepts, differentiating itself by focusing on the social aspects of learning and intelligence. Key concepts that have influenced the development of SL include:
Wisdom of the Crowds and Shared Intelligence (ShI): Documented by Galton in 1907 [40], wisdom of the crowds underscores the superiority of collective judgment over individual expertise. Its relevance in computing is demonstrated through platforms like Wikipedia, aligning with SL’s emphasis on collaborative and socially driven knowledge construction. ShI [41], which gained prominence with the growth of computer networks in the 1960s, emphasizes the distribution of intelligence through collaboration. This forms a foundational aspect of SL, where social interactions play a crucial role in the learning process.
TABLE III: Comparative Differentiation of SL from Existing Paradigms
Paradigm Core Concept SL’s Differentiation
CrI v.s. SL Aggregates many independent inputs, assuming collective wisdom leads to optimal outcomes [40]. Promotes genuine collaboration and shared learning. Interactions  in SL are processed and synthesized, not just aggregated.
ShI v.s. SL Emphasizes only distributing tasks and information [41]. Harmonizes distribution with dynamic social contexts,  making learning socially contextualized and adaptive.
SI v.s. SL Uses simple decentralized entities for collective goals [38]. Infuses decentralized model with intricate learning  algorithms, adding a social awareness component.
MAS v.s. SL Systems wherein agents engage to accomplish tasks [43]. Agents in SL not only interact but learn from these interactions, creating a richer knowledge repository.
CI v.s. SL Focuses on agent collaboration for mutual goals [44]. Integrates social dynamics, emphasizing learning, growing, and evolving through collective experiences.
Distributed Intelligence (DI) and SI: These concepts, originating from Hutchins’ theory [42] and Beni and Wang’s work [38], emphasize decentralized and collective behavior in systems. DI explores how cognition extends beyond individuals to networks, resonating with SL’s focus on collective processes. SI, inspired by natural phenomena like ant colonies, illustrates how simple rules at the individual level can lead to complex, intelligent behavior in groups.
Multi-agent Systems (MAS), Collective Intelligence (CI) and ML Ensembles: Originating from distributed AI [43], MAS encompasses autonomous agents collaborating, an idea further expanded in SL to incorporate social dynamics. CI, popularized by Lévy [44], focuses on collaborative efforts for shared objectives, mirroring SL’s emphasis on social collaboration. Introduced by Breiman [45], these methods in ML, such as random forests and boosting algorithms, parallel SL’s approach by integrating diverse models for enhanced learning, akin to SL’s focus on integrating diverse social interactions.
Figure 4: Relationships among related AI paradigms.
Crowd Intelligence: Emerging in the early 2000s, it harnesses the collective wisdom of large-scale online communities to solve complex problems [46]. It is based on the idea that the aggregated intelligence of a large group of non-experts can rival or surpass that of experts. Crowdsourcing involves distributing tasks to a large number of people via the internet.
SL’s integration of social context into AI is further detailed in Table II, contrasting it with these existing paradigms. Fig. 4 visually represents the relationships among these paradigms, highlighting SL’s innovative approach to embedding social dynamics into collective intelligence. In summary, SL represents a shift towards socialized aspects of AI, emphasizing collaborative and socially-aware problem-solving. This shift necessitates the development of sophisticated communication systems and network architectures, as discussed in our comparative analysis in Table III. Such advancements are vital for realizing the full potential of socialized AI paradigms like SL.
III EI Challenges and SL Solutions
To integrate SL into EI, we must first address inherent EI challenges such as communication efficiency, resource allocation, and data security, as shown in Fig. 5.
III-A EI Challenges Addressed by SL
Despite the advantages of EI over traditional cloud computing, it presents several challenges:
III-A1 Communication Cost
In EI, model synchronization during training, particularly with larger models and uncertain convergence, intensifies communication challenges for distributed processing and scalability. Differences in edge devices’ capabilities and data amplify these costs. Resource-constrained devices, such as smartphones, face hardware and bandwidth limitations that hinder model transmissions.
Furthermore, disparate data sizes and non-independent and identically distributed (non-IID) distributions complicate communication. Harnessing SL within EI requires addressing these communication complexities [50]. To address these complexities effectively, exploring innovative collaborative methods for efficient network designs and data transmission is essential. Such methods should align with the unique demands of EI environments, aiming to enhance model synchronization and reduce bandwidth constraints.
III-A2 Resource Allocation
EI’s transition to efficient distributed processing and local resource management presents challenges in managing a varied landscape of edge systems [9], from powerful servers to simple IoT devices, requiring strategic resource allocation. Especially when dynamic edge environments are constantly changing as devices connect and disconnect [51]. SL integration within EI can address these challenges, enhancing both efficiency and sustainability. This strategic allocation requires sophisticated algorithms capable of dynamically managing computational and communication resources, which is vital for maintaining system performance in the diverse and changing landscape of edge environments.
Figure 5: The motivation for integrating EI and SL.
III-A3 Privacy and Security
The increased exposure to diverse security protocols and edge heterogeneity necessitates enhanced privacy and security measures against multiple vulnerabilities. Data processed and stored at the edge faces increased exposure from diverse security protocols [50]. The distributed nature of EI introduces multiple vulnerabilities. Diverse threats and edge heterogeneity complicate uniform security measures.
SL integration in EI could enable adaptive security models that detect and evolve with network threats [52]. Maintaining a balance between EI’s vulnerabilities and security solutions is crucial. To achieve this balance, developing and integrating adaptive security protocols is imperative. These protocols should cater to the varied and evolving security needs of decentralized networks, ensuring data integrity and privacy across diverse edge environments.
III-B SL Solutions for EI Challenges
We explore the ways through which SL can be integrated into the EI system to address its inherent challenges. By integrating the principles of SL, we can adapt and optimize EI systems, making them more resilient, efficient, and intelligent.
III-B1 Communication Efficiency through SL
SL introduces a cooperative, adaptive approach to enhance communication efficiency in EI, addressing challenges posed by diverse edge devices and varying data distributions. Leveraging shared knowledge across devices, SL optimizes transmission, ensuring only pertinent information is exchanged. Such an intelligent, collective approach paves way for reduced communication overhead, especially in complex, non-IID scenarios [50].
III-B2 SL-driven Resource Allocation
In the intricate landscape of EI, resource management not only involves communication but also encompasses caching, computing, and control. In EI’s intricate landscape, SL aids in managing the diverse demands of computation, radio resources, and task allocation by leveraging collective knowledge to anticipate and optimize resource use. This socialized approach, backed by real-time adaptability, empowers EI systems to meet the dynamic requirements of modern tasks and applications efficiently [53].
III-B3 Enhancing Privacy and Security with SL
The decentralized structure of EI presents privacy and security challenges. SL addresses these by emphasizing knowledge sharing over direct data exchange. SL enhances privacy and security by encouraging devices to share insights and learning experiences, rather than raw data, making the network more robust against security threats. This not only safeguards individual data but also strengthens the network’s collective intelligence, protecting against data breaches and misuse [54, 52].
III-B4 Optimized Caching using SL
EI aims for efficient caching solutions to meet the growing demands for fast data access. Traditional centralized caching methods struggle to address the dynamic and diverse needs of edges. SL improves caching in EI by enabling devices to cooperatively predict and adaptively cache data based on collective behaviors, enhancing performance in real-time. Rather than working in isolation, devices share insights on frequently accessed data, intelligently populating caches and reducing unnecessary transfers to ensure pertinent data is available [55].
III-B5 SL for Effective Computation Offloading
Computation offloading in EI ensures efficient task allocation across devices. SL facilitates informed offloading decisions by allowing devices to share knowledge, enabling them to determine the most suitable device for a task’s requirements based on operational factors. The insights from SL ensure offloading strategies evolve and adapt to task and environmental changes. Hence, SL optimizes computing efficiency, contributing to energy conservation and enhanced user experience [56].
SL, with its emphasis on collaborative learning and shared experiences, offers the much-needed intelligence and adaptability to address the inherent complexities of EI. To better support EI deployment, we tailor SL to socialized architecture, training, and inference, as shown on the right side of Fig. 5. SL is detailed in Sec. IV, V, and VI.
III-C Lessons Learned
The convergence of EI and SL illuminates synergistic gains while highlighting critical challenges such as agent coordination and privacy concerns in collaborative settings. The symbiotic potential of EI and SL is evident in its capacity to alleviate communication overhead, streamline resource management and maintain privacy. SL’s approach to collaborative learning navigates through EI’s complexities, offering adaptive problem-solving and efficient communication where traditional methods may falter. This cooperative paradigm not only refines information exchange but also underscores the need for robust and adaptable communication networks and computing architectures, essential for dynamic resource allocation and processing in EI systems.
Resource optimization extends to computing and control processes, and this necessitates harnessing SL’s approach to optimize resources, leveraging shared intelligence to improve usage efficiency. SL also protects privacy by sharing insights over data, enhancing network intelligence and stopping data misuse. Beyond providing immediate solutions, the integration of SL into EI aims to create inherently adaptive systems, poised to evolve with emerging technological demands and ensure long-term relevance.
IV Socialized Architecture for EI
We establish the foundational concepts underpinning SL. Then, we delve into the social stratification that characterizes different architectural choices. In addressing these architectural choices, it’s crucial to emphasize the role of advanced communication technologies and robust network designs. These aspects are integral in ensuring efficient inter-layer communication and resource allocation, especially when dealing with complex system interconnections inherent in EI.
IV-A Functional Layers in Socialized Architecture
The EI architecture’s socialized structure includes device, edge, and cloud layers, tackling system interconnection challenges through optimized operations and resource use.
IV-A1 The Main Role in the SL Architecture
In the SL, three key layers operate collaboratively. The Device Layer includes diverse end devices like UAVs, AR tools, health sensors, IoT devices, and smart home gadgets [57, 58, 59, 60]. These devices at the Device Layer, crucial for data production and collection, process data locally or offload it based on Quality of Experience (QoE) and QoS requirements. Edge Layer provides computing and storage, mainly through ESs and gateways, orchestrating services and tasks for efficient data handling and resource optimization [61, 62]. Cloud Layer enhances traditional cloud computing with advanced capabilities for handling complex computations and large-scale storage, integrating closely with edge layer [63].
TABLE IV: Comparative Analysis in Different Aspects
Aspect Criteria Device Edge Cloud
Resource Allocation [64] Heterogeneity High Moderate Low
Algorithm
Requirement
High Moderate Low
Computing Limited Moderate High
Storage Limited Moderate Abundant
Energy
Efficiency
High Moderate Low
Data Management [65]
Local
Collection
High Moderate Low
Encryption Low Moderate High
Analytics Basic Moderate Advanced
Task Execution [66]
Real-time
Need
High High Moderate
Energy
Consideration
High Moderate Low
Complexity Simple Moderate Complex
Computing Power [67] Speed Slow Moderate Fast
Real-time
Capability
High Moderate Low
IV-A2 Comparative Analysis in Different Aspects
We analyze differences between the device, edge, and cloud layers, focusing on resource allocation, data management, task execution, and computing power to guide efficient solution design.
Resource Allocation: resource allocation differs notably among the three layers. The device layer, limited by its heterogeneity, has scarce resources, necessitating optimal algorithms for managing computation, storage, and energy [64]. While the edge layer has more resources, it’s still less capable than the cloud, emphasizing low-latency service and local data processing. The cloud layer, resource-rich, is mainly used for tasks demanding substantial computing power.
Data Management: The scope and sophistication of data management vary among the layers. The device layer centers on local data gathering with basic processing [65]. The edge layer handles advanced data tasks like preprocessing, encryption, and limited analytics. In contrast, the cloud layer is the epicenter for sophisticated data analytics and ML, providing the most extensive data solutions.
Task Execution: Task handling differs among the layers. The device layer manages simple, immediate tasks like data collection and preliminary analysis. The edge layer takes on tasks offloaded from devices, focusing on real-time actions with critical latency and energy concerns [66]. The cloud layer addresses complex tasks, such as in-depth data analysis and ML, leveraging sufficient resources without tight time limits.
Computing Power: The device, edge, and cloud layers have varied computing capabilities. The device layer is geared towards simpler tasks due to its limited power [67]. The edge layer, more powerful than devices, ensures swift processing of offloaded tasks but is still limited compared to the cloud. The cloud layer has abundant resources for intricate tasks but may lag in real-time processing because of its location. Each layer caters to distinct needs across applications.
Table IV provides a summarized comparison. To effectively cater to these distinct needs, the integration of scalable computing resources and dynamic network topologies is essential. This integration facilitates a harmonious flow of data and tasks across layers, optimizing system performance and enabling more effective management of computational and communication resources.
IV-A3 Analogies to Human Society
In analogy with human society, devices are like frontline workers handling tasks with limited resources. The edge layer, similar to local officials, focuses on local needs and policies. The cloud layer, akin to top-level leaders, offers vast resources but may face challenges due to its distance. Each layer, with its distinct role, plays a vital part in optimizing resources and cooperation, thereby enhancing the overall performance of the system.
IV-B Social Stratification in Socialized Architecture
Social stratification in SL reveals the resource and task hierarchy in computing layers. This section discusses each architecture’s strengths, challenges, and overall performance.
IV-B1 Non-Layer Architecture
Social stratification in SL begins with the basic non-layer architecture, forming a foundation for understanding more complex, centralized structures.
TABLE V: Comparison of Various Commonly used Architectures
Layer
Type
Specific
Architecture
Latency
Energy
Efficiency
Scalability
Fault
Tolerance
Application
Suitability
Advantages Limitations
Non
Centralized
[68, 69, 70]
Low High Low Low High determinism Predictable
Single point
of failure
Intra
Edge-edge
[71, 72, 73, 74]
Moderate Moderate High Moderate
Geographically
dispersed
Scalability and
resource sharing
Complex
coordination
Device-device
[18, 75, 76, 77]
High Moderate High High
High-mobility
scenarios
High scalability,
low latency
Limited
device
resources
Inter
Cloud-edge
[78]
Low to
Moderate
Moderate High Moderate
Media streaming,
IoT
Load
distribution
Limited
capabilities
Edge-device
[79, 80]
Low High Moderate Moderate
Real-time
applications
Low latency,
high energy
efficiency
Complex resource
management
Cloud-device
[81, 82]
Moderate
Moderate
to High
High Moderate Hybrid scenarios
Scalability,
flexible
resource
allocation
High latency
Cloud-edge
-device [83]
Low to
Moderate
High High High
Complex, multi-tier
applications
Highly
flexible
High management
complexity
Ad-hoc [84, 85] High
Low to
Moderate
High High
Dynamic and
unpredictable
scenarios
self-
organizing
Suboptimal
resource
allocation
IV-B1a Centralized Architecture
In contrast to other collaborative approaches, centralized strategies operate under a single governing entity, either a cloud or edge. Such a centralized system takes responsibility for resource management across all nodes within that layer. This finds particular favour when:
  • Resources are homogeneous across entities.
  • The processing capabilities of the entity are similar.
  • There is a high degree of uncertainty in predicting resource demands across individual entities.
Centralized architecture yields more efficient and reliable outcomes compared to decentralized ones. Guo et al. addressed computational offloading with a centralized heuristic algorithm in dense mobile networks [68]. Erfan et al. introduced DRF-ER for fair multi-resource allocation in EC, adapting the dominant resource fairness (DRF) model to include external resources [69]. Additionally, Kostis et al. proposed a centralized scheduler to balance queues and minimize core interference, enhancing server efficiency [70].
IV-B2 Intra-layer Architecture
Two primary architectures, namely edge-edge and device-device, have evolved as significant shifts from traditional centralized models.
IV-B2a Edge-edge Architecture
In EI, edge-edge architecture marks a shift from conventional centralized models. It uses edge interactions for better resource and workload distribution, which is beneficial for geographically dispersed networks. A key study is Huang et al.’s introduction of parked vehicle EC (PVEC), which lets vehicle EC (VEC) servers utilize parked vehicles’ (PV) idle resources for workload allocation [71]. Similarly, Na et al. used edge gateways to reduce edge workloads and maximize IoT efficiency [72]. Miao et al. also explored dynamic task offloading between edges for user QoE [73]. Thai et al. proposed a tiered collaboration model for cost-minimizing horizontal and vertical offloading [74].
IV-B2b Device-device Architecture
While edge-edge architectures have been well-investigated, the potential of device-device collaborations, especially for executing deep neural networks (DNNs), remains largely underexplored. Device-device architecture offers a paradigm shift with advantages:
  • Enhanced Local Resources: Collaboration allows for more efficient utilization of local computing power.
  • Improved Privacy: Local processing minimizes the need for external data transmission.
  • Reduced Bandwidth Dependency: Localized operations lessen the burden on network resources.
These advantages become particularly salient in high-mobility or remote environments, where traditional infrastructure might be lacking [18].
In terms of empirical studies, Zeng et al. developed CoEdge, a distributed DNN computing system that partitions workloads to adapt to device capabilities and network conditions [75]. Ranji et al. focused on improving energy efficiency and latency in device-to-device (D2D) architectures [76]. Liu et al. offered a novel cellular D2D-MEC system, highlighting effective task offloading and resource allocation strategies [77].
IV-B3 Inter-layer Architecture
Inter-layer architectures, encompassing cloud-edge, edge-device, and cloud-edge-device collaborations, leverage multi-layer resources to improve latency, energy use, and computational load.
IV-B3a Cloud-edge Architecture
The cloud-edge architecture minimizes latency and enhances user QoE, moving beyond the slower cloud-only systems. For instance, edge caching can immediately update product views in mobile shopping applications, providing faster responses to customer requests. In the domain of multimedia streaming, video transcoding also is an indispensable, yet resource-intensive operation typically relegated to clouds [78]. This approach, while effective, has the drawback of added latency. Hence, researchers like Yoon et al. suggest offloading these tasks to edge to achieve efficient and scalable solutions [86]. Similarly, studies such as those by DMRA et al. propose to expand edge layer’s role, focusing on contractual arrangements between service providers and edges, aiming to optimize resource use and minimize latency [87].
IV-B3b Edge-device Architecture
The edge-device architecture in EI focuses on minimizing latency and improving energy efficiency by leveraging closer computational resources. This approach enables data processing at the edge and on the device itself, avoiding the need for data to travel to remote clouds. Such an architecture is especially beneficial for applications where real-time performance is crucial. This architecture particularly excels in environments where rapid response times are essential for functionality or user experience. Many researchers have explored various facets of this collaborative architecture. For instance, the Edge4Sys framework suggests that the edge-device synergy can yield significant reductions in both network traffic and computational latency [79]. On a related note, advanced techniques such as ‘early exit’ have been developed to make this collaboration more efficient [80].
IV-B3c Cloud-device Architecture
The architecture emerges to alleviate latency and energy concerns associated with inference in cloud-only systems. By enabling a hybrid approach that optimizes task allocation, these architectures facilitate collaborative processing and reduce unnecessary data transmissions, thereby balancing computational needs with efficiency. Research into cloud-device architectures has led to innovations such as Hu et al.’s object detection system, which delegates complex tasks to cloud and simple ones to local devices [81], and Duan et al.’s binary search algorithm for optimal DNN partitioning [82]. Systems tend to be more effective in low-mobility settings like surveillance systems, with a focus on minimizing latency. For scenarios with high mobility, different approaches such as SPINN are necessary [88].
IV-B3d Cloud-edge-device Architecture
The cloud-edge-device framework combines the computing power of cloud computing with the edge’s low latency. This cloud-edge-device framework offers refined task and data distribution, thereby improving scalability and efficiency. While it can alleviate network bandwidth stress and reduce transmission latency, the edge’s limited computing capabilities for large-scale models can hinder its full potential. Research, such as the studies conducted by Morshed et al. analyze the complexities of developing cloud-edge-device collaborative DL algorithms [83]. These algorithms are designed to be resource and data-aware, considering a multitude of factors like heterogeneity data models and resource availability. Yet, with an increase in the number of involved entities, focus optimization shifts to overall system cost and stability.
IV-B3e Ad-hoc Architecture
Ad-hoc architecture, which allows for spontaneous resource allocation, is crucial in volatile or dynamic settings. It allows spontaneous resource allocation based on current system needs and limitations. In emergencies, resources might be allocated based on the event’s severity and available assets. Wang et al. suggest this strategy can be vital when traditional structures fall short due to fluctuating needs [84]. Ad-hoc architecture in SL offers flexibility for dynamic settings but faces challenges in efficient resource allocation, balancing responsiveness with system efficiency. Despite existing the above challenges, their adaptive nature renders them particularly relevant in certain contexts where responsiveness is paramount, like the unsupervised learning scenarios emphasized by Wang et al. [85].
Table V outlines the attributes of architectures. Centralized architectures offer control and consistency, while intra- and inter-layer ones offer flexibility for dynamic needs. The selection of an appropriate architecture should be based on the specific requirements of the EI system.
IV-C Lessons Learned
TABLE VI: Summary of Key Symbols in Section V
Symbol Description
$F(\cdot)$ Socialized learning function
$\theta,\theta_{t},\theta_{i}$ Model parameters
$N$ Number of devices or participants
$\mathcal{L},\mathcal{L}_{s},\mathcal{L}_{t}$ Loss functions
$D_{i}$ Local dataset for device $i$
$\pi,\pi^{G},\pi^{g}$ Policies in reinforcement learning
$f_{C},f_{S}$ Device and server computation functions in split learning
$P_{i}$ Personalized model for the $i$-th edge node
$w_{G},w_{i}$ Global and local model weights in FL
$Z_{i},Z_{\text{combined}}$ Feature representations and combined representation in multi-modal learning
$Q,K,V$ Query, key, and value in attention mechanism
$X,Y$ Treatment and outcome variables in causal learning
$\lambda$ Regularization weight or trade-off parameter
$\mathcal{D}$ Set of expert demonstrations in imitation learning
$H$ Cross-entropy function in knowledge distillation
$\alpha$ Weight for the distillation loss or learning rate
$P_{\text{ensemble}},P_{n}$ Ensemble and individual model predictions
$D,G$ Discriminator and generator networks in adversarial learning
$M,m_{i}$ Set of all modalities and individual modality in modality-agnostic learning
Integrating SL in EI systems accelerates distributed intelligence processing, necessitating robust encryption and advanced analytics. Key challenges include scalability to handle evolving learning algorithms, heightened security and privacy concerns, and ensuring efficient energy use. The variety of devices and the need for seamless software-hardware integration further complicate synchronized learning across platforms. Reliable real-time processing and interoperability across diverse devices and models are crucial, particularly for critical SL applications. This integration must strike a balance between centralized control and distributed resilience, optimizing resource allocation and making strategic processing decisions.
Future advancements should concentrate on refining processes to enhance scalability, interoperability, and the development of robust networking strategies. This will support the increasing complexity and scalability needs of SL in EI, ensuring robust and efficient performance in varied operational environments. Strategic leveraging of layer strengths and a proactive approach to security and resource management will be paramount.
V Socialized Training for EI
In this section, we explore the intricacies of socialized training mechanisms within the broader context of EI. Our analysis is structured around three aspects of social interaction: cooperation, submissive and conflict interaction. A summary table of the key symbols is described in Table VI.
V-A Intra-layer Cooperation in Socialized Training
V-A1 Mutual Assistance
Humans overcome various challenges through mutual assistance. Hence, devices within specific layers compensate for individual limitations through information and resource exchange, enhancing performance and robustness.
V-A1a Decentralized Learning
Decentralized learning, much like human collaboration, allows devices in the EI framework to cooperate peer-to-peer, bypassing central authorities [89], as shown in Fig. 6. Devices share and combine local datasets or partial models, thus compensating for limitations in resources or information.
The SL function Eq. (1) for decentralized learning can be represented as $F([\text{devices}],[\text{mutual assistance}],\varnothing,$ $\text{train},\text{distributed}\text{ nodes})$. The optimization objective is:
$\min_{\theta}\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(D_{i},\theta),$ (2)
where $\theta$ represents the global model parameters, $N$ is the total number of devices, $\mathcal{L}$ is the loss function, and $D_{i}$ is the local dataset for device $i$.
Figure 6: Illustration of decentralized learning in socialized training.
Decentralized learning offers computing efficiency, scalability, and data locality, with SL building on these advantages to enhance privacy and resilience by maintaining data on local devices. These methods are notably robust against challenges such as data poisoning and system bottlenecks, depending on effective decentralized averaging and network topologies [23, 90]. Popular techniques for data aggregation in such environments include consensus-based algorithms and diffusion strategies [91].
Communication Optimization in Decentralized Learning: Decentralized learning heavily relies on efficient communication among devices to achieve convergence and optimal performance. Minimizing communication overhead while ensuring fast and reliable information exchange is a critical challenge. Researchers have proposed various techniques to address this issue, such as graph-based aggregation schemes, quantization, and sparsification of model updates. These methods aim to reduce the amount of data transmitted between devices while maintaining the quality of the learned models. Moreover, the design of communication-efficient decentralized learning algorithms often involves joint optimization of learning and communication objectives, taking into account factors such as bandwidth limitations and device heterogeneity [92].
Figure 7: Illustration of distributed reinforcement learning in socialized training.
V-A1b Distributed Reinforcement Learning (DRL)
DRL facilitates coordinated decision-making in dynamic settings [93]. Especially in multi-agent reinforcement learning (MARL), devices interact with their environment for shared objectives [94], as shown in Fig. 7. Devices share critical information, like state-action pairs or local policies, to improve collective decision-making. This collaborative behavior counters challenges of large state-action spaces, delayed rewards, and heterogeneous device behavior, leveraging mutual assistance to achieve higher performance than can be attained individually.
[…]
$\max_{\pi}\mathbb{E}{\tau\sim\pi}\left[\sum_{t=0}^{T}\gamma^{t}R(s_{t},a_{t})\right],$ (3)
where $\pi$ is the policy, $\tau$ represents the trajectories, $R(s_{t},a_{t})$ is the reward function, $\gamma$ is the discount factor, and $T$ is the time horizon.
In scenarios such as autonomous driving, where centralized control is not practical, devices collaborate using local communications on a network graph in MARL, ensuring peer-to-peer interaction for mutual assistance. Zhang et al. advanced this field by introducing decentralized actor-critic algorithms with function approximation and an entropy-regularized policy gradient method that relies on neighborhood information sharing for multi-environment, multi-agent policy learning [95].
V-B Inter-layer Cooperation in Socialized Training
V-B1 Unidirectional Cooperation - Assistance
In human society, assistance involves providing resources without expecting reciprocation. Similarly, in machine society, devices offer intermediate features or perform supplementary tasks to enhance the performance of their counterparts.
V-B1a Split Learning
Split learning [96] exemplifies unidirectional assistance in inter-layer cooperation. Fig. 8 illustrates how the ML model is partitioned across multiple computational entities or ‘layers’. Devices process initial computation layers, forwarding outputs to ESs for further processing. Devices benefit from ESs’ computational capabilities. Shared features let ESs process complex tasks while preserving data privacy and reducing communication overhead. Unlike approaches like FL, which update a global model across devices, it requires less data exchange.
[…]
$\mathcal{L}=\text{Loss}\left(f_{S}\left(f_{C}(x;W_{C});W_{S}\right),y\right),$ (4)
where $\mathcal{L}$ is the total loss, $f_{C}$ and $f_{S}$ represent the device and server computations with parameters $W_{C}$ and $W_{S}$, respectively, $x$ is the input data, and $y$ is the true output.
Figure 8: Illustration of split learning in socialized training.
Split learning is suitable for scenarios requiring data privacy, like medical diagnosis [24], and in systems that require efficient use of computational resources. However, a challenge is managing the communication overhead from feature exchange in multi-device settings. The devices unilaterally assist the ESs by undertaking a part of the overall computation, thereby allowing the assisted layer to perform more complex tasks or to do so more efficiently.
Communication Efficiency in Split Learning: Split learning is a distributed learning paradigm that enables collaborative training while preserving data privacy. However, the communication overhead incurred by exchanging intermediate activations and gradients between the client and server can be significant. To mitigate this issue, researchers have explored techniques such as activation compression, selective activation sharing, and model split optimization [97]. These approaches aim to reduce the amount of data communicated between the client and server while maintaining the performance of the split learning system. Additionally, the design of communication-efficient split learning architectures may involve optimizing the placement of the split layer and the trade-off between communication cost and model accuracy.
V-B1b Edge-centralized Personalization
Edge-centralized personalization schemes feature unidirectional cooperation, where the cloud model enhances edge models for personalized tasks, as depicted in Fig. 9. The cloud model assists edges by aggregating knowledge, providing supervision, or recalibrating local models, fostering unidirectional cooperation. This approach fuses cloud computing power with edge personalization, underscoring unidirectional assistance’s potential in inter-layer machine cooperation.
[…]
$P_{i}=f(C,E_{i},D_{i};\theta),$ (5)
where $P_{i}$ is the personalized model, $C$ represents the cloud model, $E_{i}$ is the local edge model, $D_{i}$ is the local data, and $\theta$ are the governing parameters.
Figure 9: Illustration of edge-centralized personalization in socialized training.
In this context, the pioneering work by Lu et al. [98] introduces COLLA, a collaborative learning framework designed for user location prediction. In COLLA, the cloud-based model acts as a global aggregator, distilling insights from multiple edge-based models to create personalized models for each edge. The cloud-based model effectively assists edges in generating more accurate predictive models, embodying the spirit of unidirectional cooperation. Further extending this paradigm, Yao et al. [99] introduce a novel edge-cloud collaboration framework for recommendation systems that incorporates backbone-patch decomposition. This approach significantly reduces the computational burden on edges while maintaining high levels of personalization. The central model recalibrates edge backbone models to avoid local optimums through the MetaPatch mechanism.
V-B2 Bidirectional Cooperation - Cooperation
Cooperation is crucial in a machine society, where no single entity can achieve common goals independently. It depends on aligned objectives and an efficient division of labor, often through bidirectional collaboration at different layers. In training, his process entails both independent and collective modeling, characterized by continuous interaction through parameter exchanges and the potential sharing of representations. Such a process enables individuals to contribute to and benefit from the collective learning experience.
V-B2a Federated Learning
FL is a bidirectional cooperation method where participants mutually benefit from shared collective intelligence, as illustrated in Fig. 10. It allows multiple participants to collaboratively train a global model without sharing private data. In FL, a coordinating server initializes the global model and distributes it to participants. These participants then update the model using their local data and send the updated models back to the server [100]. The server aggregates these updates to refine the global model.
[…]
$w_{G}^{t+1}=\frac{1}{N}\sum_{i=1}^{N}w_{i}^{t},$ (6)
where $w_{G}^{t+1}$ is the updated global model at iteration $t+1$, $N$ is the total number of participants, and $w_{i}^{t}$ is the local model of participant $i$ at iteration $t$. Aggregation algorithms like FedAvg [101] are integral to this process, combining the local models in a way that minimizes the global loss function. This duality aligns closely with the social concept of cooperation, whereby mutual benefits are realized through collective efforts.
Figure 10: Illustration of federated learning in socialized training.
Communication Optimization in FL: FL has emerged as a promising approach for collaborative learning while preserving data privacy. However, the communication overhead associated with exchanging model updates between clients and the central server can be a significant bottleneck, especially in resource-constrained environments. To address this challenge, researchers have proposed various communication-efficient FL techniques, such as FetchSGD [102], which combines sketching and quantization to reduce communication costs, and FedMix, which approximates mixup regularization under mean augmented FL to improve model generalization and reduce communication rounds. These methods aim to strike a balance between communication efficiency and model performance in FL systems, enabling their deployment in practical settings with limited bandwidth and computational resources.
V-B2b Edge-cloud Bidirectional Collaboration
Edge-cloud Bidirectional Collaboration, as depicted in Fig. 11, capitalizes on the unique strengths of edge and cloud systems to address complex tasks such as real-time recommendations. This approach is exemplified by the Slow-Fast Learning mechanism on the Alibaba Luoxi Platform [103]. In this model, the cloud and edge systems engage in a mutually beneficial exchange, where the cloud model aids the edge model in swift predictions, and the edge model provides real-time user feedback to the cloud. The interaction between edge and cloud is shown below, highlighting essence of cooperative intelligence.
[…]
$BC=L(E,C,D,R;\theta),$ (7)
where $BC$ is the collaborative outcome, $E$ and $C$ are the states of edge and cloud models, $D$ is the real-time edge data, $R$ is the cloud’s response, and $\theta$ is the collaboration parameters.
This bidirectional collaborative framework echoes the relationship between System I and II in human cognition, as described by Daniel Kahneman et al. [104]. System I is responsible for fast, intuitive reactions, while another undertakes slower, more deliberative reasoning [105]. The coexistence and interaction between these two systems enable them to exchange prior or privileged information in a timely manner, which in turn allows for better adaptive responses to environmental demands. This collaborative approach is increasingly relevant with the rise of advanced technologies like smartphones and IoT [106].
Communication-efficient Edge-cloud Collaboration: Edge-cloud collaboration is a promising paradigm for leveraging complementary advantages of edge and cloud computing to enable intelligent and responsive services. However, the communication between edges and the cloud can introduce significant overhead and latency, hindering real-time performance of collaborative learning systems [107].
Researchers have explored communication-efficient strategies for edge-cloud collaboration, such as hierarchical compression and selective update transmission. These techniques aim to reduce the volume of data exchanged between edge and cloud nodes while preserving the quality of the learned models and the timeliness of the collaborative inference process. Optimizing edge-cloud collaboration’s communication efficiency allows SL systems to harness distributed computing resources while meeting real-world applications’ stringent latency and bandwidth constraints.
Figure 11: Illustration of edge-cloud bidirectional collaboration in socialized training.
V-B3 Bidirectional Cooperation - Exchange
In human interactions, exchange requires reciprocity and is transactional, involving the giving and receiving of resources for mutual benefit. This contrasts with cooperation, which is oriented toward joint action and shared goals. In machine societies across cloud, edge, and device layers, this principle translates into a bidirectional flow of data and model parameters, optimizing system-wide performance despite each layer’s autonomy.
V-B3a Multi-modal Learning
Multi-modal learning capitalizes on diverse data types, such as images, audio, and text, akin to human sensory processing [108]. In systems integrating cloud, edge, and device layers, each layer processes different data modalities, as depicted in Fig. 12.
[…]
$Z_{i}=f_{i}(X_{i};\theta_{i}),\quad Z_{\text{combined}}=L(Z_{1},Z_{2},\ldots,Z_{n};\Theta_{L}),$ (8)
where $Z_{i}$ is the feature representation for modality $i$, $f_{i}$ is the transformation function with parameters $\theta_{i}$, $X_{i}$ is the input data of modality $i$, $Z_{\text{combined}}$ is the combined representation, and $L$ is the fusion function with parameters $\Theta_{L}$.
One noteworthy application that captures this essence of exchange is the multi-modal federated collaborative training framework by Qayyum et al. [109]. The framework builds a composite model for diagnosing COVID-19 by harmoniously integrating two modalities of data-X-ray and ultrasound images, across different centers in a cloud-edge environment. Here, the interaction is not merely cooperative but fundamentally reciprocal, as each layer significantly contributes to the shared model and benefits from it for its own objectives.
We provide a comprehensive exploration of cooperation dynamics within ML layers, drawing parallels to human social interactions. A comparative overview is presented in Table VII, highlighting their distinct characteristics.
Figure 12: Illustration of multi-modal learning in socialized training.
TABLE VII: Comparative Analysis of Technical Approaches Reflecting Cooperation Relationships
Cooperation
Type
Technique
Social
Relationship
Social
Characteristics
Model
Interactivity
Privacy
Maintenance
Scalability Resilience
Intra-layer
Decentralized
Learning [89, 23, 90, 91, 92]
Mutual Assistance
Peer-to-peer
cooperation
High High High High
Distributed
Reinforcement Learning [95]
Coordinated
decision-making
High Medium High Medium
Inter-layer
Model Split
Learning [96, 97]
Assistance
Unilateral
assistance
Medium High Medium Medium
Edge-centralized
Personalization [98, 99]
Cloud aiding edges Medium High High Medium
Federated
Learning [100, 101, 102]
Cooperation
Reciprocal benefit and
contribute equally
High High High High
Edge-cloud Bidirectional
Collaboration [103, 104, 105, 106, 107]
Interdependent and
reciprocal
High Moderate High High
Multi-modal
Learning [108, 109]
Exchange
Exchanging diverse
data modalities
High Moderate High Moderate
V-C Intra-layer Submissive Interaction in Socialized Training
V-C1 Conformity
In society, individuals often conform to majority norms due to various pressures or desires. In machine society, individuals mine separable discriminant features or similar parameter relationships among individuals in the training process and reduce model bias by increasing the impact of discriminant features on models or maximizing the retention of similar parameter relationships between individuals.
V-C1a Causal Learning
Causal learning in EI is crucial for maintaining fairness [110] and reducing bias, particularly by facilitating universal model applicability through on-cloud training and mitigating overfitting in edge training. This approach is illustrated in Fig. 13. A fundamental concept in causal learning is estimating the average causal effect (ACE) of one variable on another. Given a set of variables, we often are interested in the causal effect of one variable $X$ (treatment) on another $Y$ (outcome). The SL function Eq. (1) for causal learning is $F([\text{servers}],[\text{conformity}],\varnothing,\text{train},\text{variables})$. The average causal effect (ACE) is:
$ACE=\mathbb{E}[Y_{1}-Y_{0}]=\mathbb{E}[Y_{1}]-\mathbb{E}[Y_{0}],$ (9)
where $Y_{1}$ and $Y_{0}$ are the potential outcomes under treatment ($X=1$) and control conditions ($X=0$), respectively, and $\mathbb{E}[\cdot]$ denotes the expectation.
Figure 13: Illustration of causal learning in socialized training.
Integrating model compression with causal learning for enhanced generalization is demonstrated in [111]. Several other research studies have extensively explored the convergence of causality and out-of-domain generalization [112, 113, 114, 115, 116, 117]. For instance, [114] underscores the importance of preserving semantics intrinsic to the target domain by leveraging disentangled causal mechanisms. The study in [116] posits the enduring relationship between causal features and classifications across domains and champions the Markov Blanket [118] for strategic causal feature selection. The proposal by [115] introduces a causal regularizer, aiming to resurrect the causation between predictors and outcome variables when faced with unpredictable distributions.
V-C1b Attention Mechanism
Such a mechanism, by selectively amplifying certain portions of the input while downplaying others, embodies the principle of conforming to pivotal information [119]. Architectures like the Transformer, employing self-attention, underscore this conformity by weighting the importance of each word in relation to others. This evolution in design philosophy has ushered in models such as BERT [120] and GPT [121], which further epitomize ‘conformity’ by foregrounding contextually salient words in expansive text corpora. This mechanism is illustrated in Fig. 14. The SL function Eq. (1) for the attention mechanism is $F([\text{servers}],[\text{conformity}],\varnothing,\text{train},\text{data})$. The attention mechanism is defined as follows:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (10)
where $Q$, $K$, and $V$ are the query, key, and value matrices, respectively, and $d_{k}$ is the dimension of the key vectors. $QK^{T}$ calculates attention scores, which, after normalization and conversion by the Softmax function, help the model determine focus areas in the input data.
Figure 14: Illustration of attention mechanism in socialized training.
The attention mechanism reflects social conformity by emphasizing prominent familiar regions of an image. Both channel and spatial attention mechanisms echo the varied nuances of social conformity. Channel attention mirrors our focus on prevailing trends, distinguishing vital channels akin to distinctive social figures [122]. Similarly, spatial attention resonates with the social spotlight on specific locations or events [123]. This selective focus of attention mechanisms, evident in models like BERT, serves as an embodiment of social conformity [124]. However, it’s essential to recognize that while attention provides significant insights, it’s not an exhaustive explanation of model behavior, much like adhering to a norm doesn’t fully capture society’s complexity [125].
V-D Inter-layer Submissive Interaction in Socialized Training
V-D1 Unidirectional Submissive Interaction - Suggestion&imitation
In the same way that suggestion and imitation play roles in human interactions, with individuals adapting behaviors observed in others, models also rely on similar principles during training. Complex models act as authoritative guides, transferring knowledge to newer models through intermediate features or outcomes. New models then aim to closely replicate these features and outcomes in an attempt to ‘imitate’ the performance of more sophisticated models.
V-D1a Transfer Learning
[…]
$\quad\quad\underset{\theta_{t}}{\text{min}}\;\mathcal{L}_{t}(f(X_{t};\theta_{t}),Y_{t})+\lambda\mathcal{R}(\theta_{t}),$
where $\mathcal{L}_{s}$ and $\mathcal{L}_{t}$ are the loss functions for the source and target domains, $f$ is the model with parameters $\theta$ and $\theta_{t}$, $X_{s}$ and $X_{t}$ are the source and target data, $Y_{s}$ and $Y_{t}$ are the corresponding labels, $\mathcal{R}$ is the regularization term, and $\lambda$ is the regularization weight. $\text{Transfer}(\cdot)$ denotes the knowledge transfer function.
Figure 15: Illustration of transfer learning in socialized training.
The diversity in data distribution across edges and the cloud highlights the significance of heterogeneous transfer learning, as demonstrated in [126]. This learning approach adapts to varying data characteristics, employing symmetric transformations for domain-agnostic representations, as discussed in [127], and asymmetric transformations for domain-specific alignment, highlighted in [128]. This mirrors the social dynamic where one domain or group modulates its behavior to resemble another, often seen as superior. This concept is further exemplified by semi-supervised kernel matching domain adaptation [129], showcasing models’ capacity to assimilate and integrate more advanced insights.
V-D1b Imitation Learning
In social constructs, people often follow the guidance of authoritative figures. Similarly, in imitation learning, a novice model observes and imitates the expert model’s behavior, akin to learning through observation in social interactions. This method involves recording the expert’s actions and corresponding states, then using this information to predict actions in similar situations [130].
[…]
where $\mathcal{D}$ is the set of expert demonstrations, $\pi_{\theta}$ is the learned policy with parameters $\theta$, $\mathcal{L}$ is the loss function, and $a_{\text{pred}}$ is the predicted action for a new state $s_{\text{new}}$.
Imitation learning enables quick knowledge acquisition and application to various domains, effectively skipping learning steps. However, it depends on access to an expert model and can lead to overfitting issues [130]. Despite these challenges, imitation learning is a key technique in AI, particularly useful in fields such as autonomous driving and game AI.
V-D1c Knowledge Distillation (KD)
[…]
$\mathcal{L}_{KD}=H(y,q^{\text{student}})+\alpha H(q^{\text{teacher}},q^{\text{student}}),$ (13)
where $H$ is the cross-entropy function, $y$ is the true label, $q^{\text{student}}$ and $q^{\text{teacher}}$ are the softened outputs of the student and teacher models, and $\alpha$ is the weight for the distillation loss.
Figure 16: Illustration of knowledge distillation in socialized training.
Submissive interaction in ML prominently manifests in three KD paradigms, each drawing parallels to social interactions. Response-based KD imitates the direct responses of expert models, using these soft targets to guide student models. Feature-based KD goes further, teaching student models to mimic intermediate features of teacher models, with FitNet as a prime example [132]. It aligns student and teacher feature maps for deeper knowledge transfer. Lastly, Relation-based KD explores the relationships between layers, akin to the interplay of skills and concepts in social settings. The feature space projection matrix, capturing feature map pair relationships, is a notable method in this context [133]. Together, these KD methods reflect the ML trends from novice to expert, mirroring human learning processes.
Communication-aware Knowledge Distillation: KD is a technique for transferring knowledge from a large, complex teacher model to a smaller, simpler student model, which can be deployed on resource-constrained devices for efficient inference. However, in a distributed setting, the communication cost of transferring the teacher model’s knowledge to multiple student models can be prohibitive. Researchers have proposed communication-aware KD techniques [134], which combine FL with KD to enable collaborative and communication-efficient knowledge transfer across multiple clients. These approaches aim to minimize the communication overhead of KD while ensuring that the student models can effectively absorb the knowledge of the teacher model and achieve comparable performance on the target task. By incorporating communication-awareness into the KD process, SL systems can facilitate the efficient deployment of powerful AI models on resource-constrained devices, enabling intelligent and responsive services at the edge.
V-D1d Meta Learning
Meta learning, which aims to improve learning processes by utilizing experiences from past tasks, differs from transfer learning’s knowledge application, as shown in Fig. 17. Advances such as MELO [135], which extends model-agnostic meta-Learning (MAML) and applications involving spiking neural networks, demonstrate the adaptation of meta learning in EI. This evolution, when combined with model compression, equips models for the unique demands of EC [136].
Figure 17: Illustration of meta learning in socialized training.
Furthermore, the challenges posed by limited datasets in traditional ML methods echo the real-world scenarios where apprentices may only have a handful of opportunities to observe and learn from a master. Just as humans intuitively learn to learn, meta learning thrives on encapsulating patterns, problem properties, and past performances to accelerate the learning of new tasks, even with sparse data. In doing so, it not only epitomizes the spirit of human adaptability but also reinforces the social essence of learning through suggestion and imitation, achieving performance levels that potentially outstrip conventional methods.
[…]
$\theta^{*}=\arg\min_{\theta}\mathcal{L}_{\text{meta-train}}(\theta),\quad\theta_{i}^{\prime}=\theta-\alpha\nabla_{\theta}\mathcal{L}_{\text{train}}(f_{\theta}(x_{i}),y_{i}),$ (14)
$\mathcal{L}_{\text{meta-test}}(\theta^{\prime})=\mathcal{L}_{\text{test}}(f_{\theta^{\prime}}(x^{\prime}),y^{\prime}),$
where $\theta^{*}$ are the optimal parameters learned during meta-training, $\mathcal{L}_{\text{meta-train}}$ and $\mathcal{L}_{\text{meta-test}}$ are the meta-training and meta-testing loss functions, $f_{\theta}$ is the model with parameters $\theta$, $(x_{i},y_{i})$ and $(x^{\prime},y^{\prime})$ are the training and testing data pairs, $\alpha$ is the learning rate, and $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ are the loss functions for the training and testing tasks.
V-D2 Bidirectional Submissive Interaction
In machine societies, entities participate in bidirectional submissive interactions, meaning they both acquire and contribute knowledge. This dynamic allows them to switch roles between learner and mentor, ensuring that individual insights lead to collective system improvement and optimize overall efficiency.
TABLE VIII: Comparative Analysis of Technical Approaches Reflecting Submissive Interactions
Submissive
Type
Technique
Social
Relationship
Social
Characteristics
Model
Interactivity
Complexity Flexibility Scalability Resilience
Intra-layer
Causal
Inference
[110, 111, 112, 113, 114, 115, 116, 117, 118]
Conformity
Conforming to reasons
or norms
Low High Moderate Moderate High
Attention
Mechanism
[119, 120, 121, 122, 123, 124, 125]
Focusing on dominant
features
Low Moderate High High Moderate
Inter-layer
Transfer
Learning
[126, 127, 128, 129]
Suggestion & Imitation
Drawing guidance from
experienced figures
Moderate Moderate Moderate High Moderate
Imitation
Learning [130]
Directly mimicking
the actions of an expert
Moderate Low Low Moderate Low
KD
[131, 132, 133, 134]
Learning from an expert,
inheriting wisdom
Low Moderate High High High
Meta-
Learning
[135, 136]
Learning the nuances of
a craft from experts
High High Very High High High
DA
[137, 138, 139, 140, 141, 142, 143, 144]
Bidirectional Submissive
Mutual adaptation
between entities
High High High Moderate High
MTL
[145, 146, 147, 148]
Collaborative learning
with interdependent tasks
Very High High Very High Very High Moderate
Figure 18: Illustration of domain adaptation learning in socialized training.
V-D2a Domain Adaptation (DA) Learning
DA functions in a bidirectional or even multi-directional realm of transfer learning [137]. It leverages knowledge from a source domain to adapt to the target domain’s data distribution. This exchange is depicted in Fig. 18. Instance-based DA [138, 139] reweights source instances to bridge disparities between source and target samples. Feature-based DA [140, 141, 142] creates a unified space for the cohesion of source and target features. Gani et al.’s approach [142] utilizes generative adversarial networks (GANs) to create consistent features across domains.
[…]
$\min_{\theta}\mathcal{L}(\theta)=\mathcal{L}_{s}(\theta)+\lambda\cdot\text{MMD}^{2}(H_{s},H_{t}),$ (15)
where $\mathcal{L}_{s}$ is the source domain loss, MMD is the maximum mean discrepancy, $H_{s}$ and $H_{t}$ are feature representations for the source and target domains, and $\lambda$ is the trade-off parameter.
Figure 19: Illustration of multi-task learning in socialized training.
Insights from the target domain enhance the source’s knowledge. Liu et al. [143] describe this with their cycle self-training (CST) model, which facilitates knowledge exchange between domains. Such intricate dynamics of DA become indispensable in realms like EI where data distributions diversify across devices. Representative of this is the MobileDA framework by Yang et al. [144], which not only combats domain shifts in EI devices but also facilitates seamless knowledge exchange between centralized servers and edge entities.
V-D2b Multi-task Learning (MTL)
MTL [145] orchestrates bidirectional submissive interactions among diverse tasks, as shown in Fig. 19. Just as diverse domains in society, such as the realms of culture, economy, and governance, simultaneously influence and adapt to each other, MTL undertakes the collaborative act of sharing model parameters and feature representations.
[…]
$\mathcal{L}_{MTL}=\sum_{t=1}^{T}\alpha_{t}\mathcal{L}_{t},$ (16)
where $\mathcal{L}_{MTL}$ is the total multi-task loss, $\mathcal{L}_{t}$ is the loss for task $t$, $\alpha_{t}$ is weight for task $t$, and $T$ is total number of tasks.
Within the MTL, the categorization into hard sharing [146] and soft sharing [147] mirrors social dynamics where certain norms are universally rigid (hard sharing) while others permit regional variances (soft sharing). Maurer et al.’s exploration illustrates how shared features, envisioned as transformations through multi-layered networks [148], parallel the way societies adapt universal principles into specific regional customs. In the larger picture, MTL champions a symbiotic feature interaction across multiple tiers, harmonizing model structures and fostering task interdependencies. The result is enhanced model robustness, adaptability, and commendable prowess to tackle a plethora of complex tasks. Bidirectional submissive interactions, illustrated by DA and MTL, underscore the importance of knowledge exchange. For clarity, Table VIII provides a comparative analysis of these methods.
V-E Intra-layer Conflict Interaction in Socialized Training
V-E1 Competition
Competition is a driving force in many systems, pushing entities to strive for supremacy in various domains. In the machine society, algorithms and models mirror this competitiveness, aiming to outperform peers through constant fine-tuning, adaptation, and evolution. Each model desires to set benchmarks for others to surpass.
Figure 20: Illustration of ensemble learning in socialized training.
V-E1a Ensemble Learning
Ensemble learning epitomizes intra-layer competition where models vie to showcase their predictive prowess on a common dataset. While individual models are complete in themselves, their competitive advantages become more pronounced when juxtaposed with peers, pushing each model to highlight its strengths and compensate for its weaknesses. This concept is illustrated in Fig. 20.
[…]
$P_{\text{ensemble}}=\sum_{n=1}^{N}w_{n}P_{n},$ (17)
where $P_{\text{ensemble}}$ is the ensemble prediction, $P_{n}$ is the prediction of the $n$-th model, $w_{n}$ is the weight for the $n$-th model, and $N$ is the total number of models.
Ensemble learning employs various strategies to harness this competition. In particular, approaches like voting [149] by pooling predictions, whereas weighted voting [150] confers preferential weights based on a model’s relative performance. Further, the model averaging method [151] seeks a consensus by averaging individual model outputs. Bai et al. [152] proposed an Automated DNN ensemble (AES), which selects optimal DNN models and tunes the ensemble according to nuances of the input data. Such intricate competitive dynamics have found success across many applications, from fraud detection [153] and medical diagnosis [154] to image recognition [155], emphasizing its effectiveness in harnessing intra-layer competition for superior predictive outcomes.
V-F Inter-layer Conflict Interaction in Socialized Training
V-F1 Unidirectional Competition - Compulsory
Compulsory interaction is characterized by the dominance of one group and the subsequent compliance of another. Often, this stems from structural hierarchies, such as in workplaces where managers direct employees or in social constructs where certain groups have influence over others. In ML, this interaction can be observed in hierarchical structures where higher-level algorithms or protocols dictate the functioning or constraints of those at lower levels. By understanding this unidirectional competition, researchers can harness the efficiency and directionality it offers while being mindful of the inherent constraints and potential lack of flexibility it introduces.
V-F1a Hierarchical Reinforcement Learning (HRL)
HRL mirrors unidirectional compulsory interactions where a dominant group or layer prescribes rules or guidelines for another group, mandating compliance. This structure is visually represented in Fig. 21. The essence of compulsory is exhibited as the superiors set overarching goals, while the subordinate layers are obligated to adhere. Dayan et al. [156] offer a quintessential example with a system partitioned into managers and sub-managers. Governed by principles like reward hiding and information hiding, this structure encapsulates unidirectional competition, with each layer, while autonomous to some extent, invariably remaining in contact with its superiors.
Figure 21: Illustration of hierarchical reinforcement learning in socialized training.
[…]
$H(s,\pi^{G},\pi^{g},\text{Transit},R)=[\text{Transit}(s,\pi^{g}(s,\pi^{G}(s))),$ (18)
$R(s,\pi^{g}(s,\pi^{G}(s)),\text{Transit}(s,\pi^{g}(s,\pi^{G}(s))))],$
where $H$ is the main function, $s$ is the state, $\pi^{G}$ and $\pi^{g}$ are the high-level and low-level policies, $\text{Transit}(\cdot)$ is the transition function, and $R$ is the reward function. Such hierarchical frameworks not only streamline learning but also afford the assimilation of prior knowledge. As in social structures, HRL embodies the directives of one group to the behaviour of another. The directives from higher layers, while potentially restrictive, shape the trajectory of the entire system, ensuring alignment with overarching objectives [157].
V-F2 Bidirectional Competition - Struggle
The concept of struggle in social dynamics, when entities compete for shared resources, power, or influence, is reflected in adversarial learning techniques. In these techniques, two components challenge each other, and it’s through this continuous push and pull that the model evolves, yielding improved outcomes.
TABLE IX: Comparative Analysis of Technical Approaches Reflecting Conflict Interactions
Conflict
Type
Technique
Social
Relationship
Social
Characteristics
Model
Interactivity
Complexity Scalability Adaptability
Outcome
Diversity
Intra-layer
Ensemble
Learning
[149, 150, 151, 152, 153, 154, 155]
Competition
Collaborative rivalry
among peers
Moderate Moderate High Moderate High
Inter-layer HRL [156, 157] Compulsory
Top-down directive
from superiors
Low High Moderate Low Moderate
Adversarial
Learning
[158, 159, 160, 161, 162, 163, 164]
Struggle
Mutual power
struggle between
opponents
High High Moderate High High
Modality Agnostic
Learning
[165, 119]
Adaptability across
diverse environments
Low High High
Very
High
Very
High
Figure 22: Illustration of adversarial learning in socialized training.
V-F2a Adversarial Learning
In adversarial learning, particularly notable in GANs [158], there is a struggle between two components: the generator and the discriminator, as illustrated in Fig. 22. The generator crafts authentic samples to deceive the discriminator. Conversely, the discriminator endeavors to discern real data from the counterfeits created by the generator. However, challenges like instability in the training process and homogeneous sample generation persist. Original GANs often produced uncontrollable image variations from random noise.
[…]
$\min_{G}\max_{D}V(D,G)=\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D(x)]+$ (19)
$\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))],$
where $G$ and $D$ are the generator and discriminator networks, $p_{\text{data}}$ is the real data distribution, $p_{z}$ is the noise distribution, and $V(D,G)$ is the value function.
Solutions emerged in the form of conditional GANs [159], introducing conditional data and Deep Convolution GCNs (DCGANs) [160], swapping fully connected layers for convolutional ones. The quest for stability birthed the Wasserstein GAN [161] with its novel loss function. Drawing parallels to social struggle, adversarial learning represents inter-layer conflicts. This dynamic framework extends beyond image generation, spanning areas like RL [162] and natural language processing (NLP) [163].
Communication-aware Adversarial Learning: Adversarial learning techniques, such as GANs, have shown remarkable success in various applications. However, in a distributed setting, the communication between the generator and discriminator networks can be costly and may impact the overall performance. Communication-aware adversarial learning techniques have been proposed to address this issue by reducing communication overhead while maintaining the quality of the generated samples. These methods may involve strategies such as gradient compression, selective parameter sharing, and decentralized architectures for adversarial learning. The design of communication-efficient adversarial learning systems often requires considering the trade-off between communication cost and the stability and convergence of the training process [164].
V-F2b Modality Agnostic Learning
[…]
$\begin{cases}x\in M=\{m_{1},m_{2},\ldots,m_{n}\},\\ y=f(x;\theta),\end{cases}\vspace{-0.5em}$ (20)
where $M$ is the set of all modalities, $x$ is the input data from any modality in $M$, $f$ is the modality-agnostic model with parameters $\theta$, and $y$ is the model output.
TABLE X: Comparative Analysis of ML Methods in Socialized Training for SL
ML Method Scope of Application Advantages Disadvantages
Decentralized Learning [89, 23, 90, 91, 92] Peer-to-peer collaboration among devices with limited resources. Efficient resource utilization, improved privacy, and robustness against single points of failure. Increased communication overhead, the potential for slower convergence.
DRL [93, 94, 95] Multi-agent systems with shared objectives and dynamic environments. Enables coordinated decision-making and adaptation to complex environments. Challenges in managing large state-action spaces, delayed rewards, and heterogeneous agents.
Split Learning [96, 24, 97] Systems with privacy constraints and heterogeneous computational capabilities. Preserves data privacy, reduces communication overhead, and enables collaboration among devices with varying resources. Potential for increased computational complexity and challenges in managing multiple local models.
Edge-centralized Personalization [98, 99] Applications requiring personalized models at the edge with centralized control. Enables personalized edge models while leveraging cloud resources and global knowledge. Increased dependency on the cloud, potential for higher latency, and privacy concerns.
FL [100, 101, 102] Scenarios with distributed data and privacy requirements. Enables collaborative learning without sharing raw data, reduces communication costs, and preserves privacy. Challenges in managing non-IID data, model divergence, and potential security vulnerabilities.
Edge-cloud Bidirectional Collaboration [103, 104, 105, 106, 107] Applications requiring real-time insights and global knowledge. Leverages the strengths of both edge and cloud computing, enables fast local decisions and global model updates. Increased complexity in managing the collaboration, potential for higher communication costs.
Multi-modal Learning [108, 109] Systems with heterogeneous data types and modalities. Enables learning from diverse data sources, improves model generalization and robustness. Challenges in data alignment, fusion, and increased computational complexity.
Causal Learning [110, 111, 112, 113, 114, 115, 116, 117, 118] Scenarios requiring unbiased and interpretable models. Enables the discovery of causal relationships, improves model interpretability and fairness. Requires domain knowledge for causal graph construction, potential for increased computational complexity.
Attention Mechanism [119, 120, 121, 122, 123, 124, 125] Applications with sequential or spatial dependencies. Enables focusing on relevant information, improves model interpretability and performance. Increased computational complexity, potential for overemphasizing certain features.
Transfer Learning [126, 127, 128, 129] Scenarios with limited labeled data or related source domains. Enables knowledge transfer from related tasks, reduces the need for labeled data, and improves model performance. Challenges in identifying suitable source domains, potential for negative transfer.
Imitation Learning [130] Applications with expert demonstrations or optimal behaviors. Enables learning from expert demonstrations, reduces the need for explicit reward functions. Requires access to expert demonstrations, potential for suboptimal policies and lack of generalization.
KD [131, 132, 133, 134] Scenarios with resource-constrained devices and pre-trained large models. Enables transfer of knowledge from large to small models, reduces computational requirements. Potential for loss of information during distillation, challenges in managing the trade-off between model size and performance.
Meta Learning [135, 136] Applications requiring fast adaptation to new tasks or environments. Enables rapid learning and adaptation to new tasks, improves model generalization. Increased computational complexity, challenges in defining appropriate meta learning objectives.
DA Learning [137, 138, 139, 140, 141, 142, 143, 144] Scenarios with domain shift or varying data distributions. Enables knowledge transfer across different domains, improves model performance and robustness. Challenges in identifying and aligning relevant domain features, potential for negative transfer.
MTL [145, 146, 147, 148] Applications with related tasks or shared representations. Enables learning from multiple tasks simultaneously, improves model generalization and efficiency. Challenges in balancing task-specific and shared representations, potential for task interference.
Ensemble Learning [149, 150, 151, 152, 153, 154, 155, 45] Scenarios requiring robust and accurate predictions. Improves model performance and robustness, reduces overfitting, and enables uncertainty quantification. Increased computational complexity, potential for increased model size and inference time.
HRL [156, 157] Applications with hierarchical task structures or long-term goals. Enables learning at different levels of abstraction, improves sample efficiency and transfer learning. Challenges in defining appropriate hierarchies, potential for suboptimal policies due to hierarchical constraints.
Adversarial Learning [158, 159, 160, 161, 162, 163, 164] Scenarios requiring realistic data generation or robust models. Enables the generation of realistic data, improves model robustness and performance. Challenges in training stability, mode collapse, and potential for adversarial attacks.
Modality Agnostic Learning [165, 119] Applications with diverse and evolving data modalities. Enables learning from various data types, improves model flexibility and adaptability. Increased computational complexity, challenges in handling the heterogeneity of data modalities.
Central to this evolution is the perceptron model [165], which drives the development of modality-agnostic learning. This model finds its roots in the Transformers networks [119], which contrary to their traditional domain of sequence transduction, have demonstrated prowess in computer vision tasks, all while consuming lesser computational resources. This dynamicity not only facilitates broader participation in EI but also sets the stage for more robust and reliable network optimization tasks.
In summary, we explore the dynamics of conflict interactions within and between layers of ML techniques. Each approach contributes uniquely to our understanding of interactivity, complexity, scalability, adaptability, and outcome diversity in ML models, as outlined in Table IX.
To provide a comprehensive overview of the ML methods discussed in this section, we present a comparative analysis of their scope of application, advantages, and disadvantages in Table X. This table enables readers to quickly grasp the key characteristics and potential use cases of each method in the context of socialized training and EI.
V-G Lessons Learned
Our examination within the context of SL and EI underscores the merging of complex human interactions and ML training. This fusion points to an emerging trend of AI models that increasingly incorporate human social behavior. The trend towards adaptable AI, influenced by frameworks like the perceiver model, mirrors social adaptability during transitions. ML’s pursuit of devising versatile algorithms reflects the need for a balance between efficiency and adaptability.
Additionally, our study delves into the dichotomy of individual versus collective efficiency, as reflected in ensemble learning’s portrayal of social individualism and collectivism. This also relates to the HRL’s directive versus autonomy debate, reflected in the balance of control and independence seen in society. Developing ML models that balance resilience and adaptability while retaining specialization is challenging. The apparent synergy between ML and social dynamics signals a future maturity for cross-disciplinary cooperation.
Figure 23: Illustration of edge-edge collaborative inference.
VI Socialized Inference for EI
This section delves into the integration of socialized inference in EI. It examines how social dynamics can improve decision-making, highlighting social interactions and their effects on inference.
VI-A Intra-layer Cooperation in Socialized Inference
VI-A1 Mutual Assistance
In EI, mutual assistance is essential for collaborative inference. Despite limitations, each device contributes unique abilities for the group’s benefit.
VI-A1a Edge-edge Collaborative Inference
As shown in Fig. 23, EI employs edge cooperation in distributed inference, leveraging collective resources for better inference. This approach negates central oversight as edges converge on shared goals. Optimization goals categorize this collaboration into:
  • Total Inference Latency Minimization: Edges accelerate inferences collaboratively, avoiding central delays. Just as members of society possess varied abilities, edge resources also are different. Addressing this involves task allocation and workload balancing. Studies such as Wan et al. [166] and Mohammed et al. [167] delve into this, proposing DL-based strategies that maximize resource utilization in EI.
  • Total Cost Minimization: Beyond speed, holistic costs, such as energy and communication overheads, reflect sustainable social cooperation goals. The focus here is on achieving a balance and distributing tasks in a manner that minimizes collective costs. Studies like those proposed by Fan et al. [168] have ventured into this area, optimizing edge collaborations for minimum costs.
[…]
s.t. $\quad C(t_{j},e_{i})\leq R(e_{i}),\quad\sum_{j=1}^{m}C(t_{j},e_{i})\leq\omega,\quad\forall i,j,$
where $E=\{e_{1},e_{2},\ldots,e_{n}\}$ is the set of edges, $T=\{t_{1},t_{2},\ldots,t_{m}\}$ is the set of tasks, $R(e_{i})$ is the available resources on edge $e_{i}$, $C(t_{j},e_{i})$ is the computational cost of task $t_{j}$ on edge $e_{i}$, $D(e_{i},e_{k})$ is the communication cost between edges $e_{i}$ and $e_{k}$, $L(t_{j},e_{i})$ is the latency of task $t_{j}$ on edge $e_{i}$, $\lambda$ is the trade-off parameter, and $\omega$ is the maximum workload allowed on each edge.
Figure 24: Illustration of device-device collaborative inference.
Edge-edge collaborative inference, seen through the lens of mutual assistance, boasts several merits: 1) Privacy and Security. The communal edge approach safeguards sensitive data, echoing the social emphasis on individual privacy. 2) Scalability and Flexibility. Reflecting the adaptability of social cooperation, this method is inherently scalable and accommodative of diverse edges. 3) Low Latency. Emphasizing timely responses, the method caters to real-time applications, paralleling societies’ emphasis on immediacy in collaboration.
While the virtues are evident, the challenges in edge-edge cooperation also mirror those in social mutual assistance: 1) Communication Overhead. Collaboration invariably involves communication, which can be burdensome, more so with increasing collaborators. 2) Heterogeneity. Just as individuals in a society vary, so do edges in capability, posing optimization challenges. 3) Fault Tolerance. Societies must contend with unpredictable challenges; similarly, edge collaboration needs to account for device failures and network anomalies.
Communication Optimization in Edge-edge Collaborative Inference: Edge-edge collaborative inference heavily relies on efficient communication among edges to achieve optimal performance and minimize latency. One key challenge is to optimize the communication network topology and routing strategies to ensure fast and reliable information exchange. This can be formulated as a joint inference and communication optimization problem, aiming to minimize total inference latency while satisfying communication bandwidth and reliability constraints. Techniques such as network encoding and cooperative communication can be applied to improve the communication efficiency and robustness in edge-edge collaborative inference scenarios [169].
VI-A1b Device-device Collaborative Inference
Device-device collaborative inference in EI highlights the combined efforts of devices, as illustrated in Fig. 24. Just as individuals in a society pool their strengths and resources to overcome challenges that are beyond their capability alone, devices, particularly those with limited computational capacities. They can collaborate, exchange information, and process data jointly.
  • Total Inference Latency Minimization: This aspect involves the strategic sharing of computational tasks, akin to public resource allocation, aimed at reducing overall inference delays. Especially critical for UAVs and other latency-sensitive applications, this strategy involves optimizing task allocation across devices, even when faced with inherent resource limitations [170]. The objective remains to prevent resource overloading on a single device, ensuring harmonious operations [171].
  • Total Cost Minimization: The collaborative EI also extends to encompass cost-effectiveness, where devices, particularly those limited by battery life and storage, operate under a paradigm of energy efficiency. This approach necessitates a balance between energy consumption for inference operations and communication, as highlighted in studies by Yang et al. [172] and Zeng et al. [173].
The objective is to optimally allocate tasks to devices to minimize both inference latency and energy consumption. The SL function Eq. (1) for device-device collaborative inference is $F([\text{devices}],[\text{mutual assistance}],\varnothing,\text{inference},\text{tasks})$. The optimization objective is:
$\underset{x_{ij}}{\text{min}}$ $\sum_{i=1}^{n}\sum_{j=1}^{m}(L_{ij}x_{ij}+E_{ij}x_{ij}),$ (22)
s.t. $\quad\sum_{i=1}^{n}x_{ij}=1,\quad\sum_{j=1}^{m}R_{ij}x_{ij}\leq R_{i},\ x_{ij}\in\{0,1\},\ \forall i,j,$
where $L_{ij}$ is the inference latency of task $t_{j}$ on device $d_{i}$, $E_{ij}$ is the energy consumption of task $t_{j}$ on device $d_{i}$, $x_{ij}$ is the assignment variable indicating whether task $t_{j}$ is assigned to device $d_{i}$, $R_{ij}$ is the resource requirement of task $t_{j}$ on device $d_{i}$, and $R_{i}$ is the available resources on device $d_{i}$.
Figure 25: Illustration of edge-device collaborative inference.
The advantages of this collaborative inference model include: 1) System independence: It promotes operational autonomy of devices, mitigating reliance on central infrastructures, and thereby reducing associated costs. 2) Reduced latency: Collaborative efforts enhance decision-making speeds, a feature reminiscent of quick social problem-solving dynamics. 3) Adaptability: The system exhibits resilience against disruptions, reflective of social robustness.
However, it also introduces challenges: 1) Privacy concerns: Increased inter-device communication potentially elevates the risk of data leakage, a dilemma paralleling privacy concerns in human societies. 2) Synchronization among heterogeneous devices: Variances in devices may complicate synchronization, an issue analogous to aligning individuals with diverse abilities in a society. 3) Task adaption: The necessity for rapid adaptability in task management, considering variables like energy and computation, reflects the dynamism inherent in social collaborations.
VI-B Inter-layer Cooperation in Socialized Inference
VI-B1 Unidirectional Cooperation - Assistance
In EI during inference, certain computing layers provide insights or intermediary results to speed up others’ processes. The aided entities gain the advantages while the helpers play a supporting role. Unlike training, where knowledge transfer is paramount, in EI’s inference phase, the focus shifts to utilizing specific layers or devices to minimize latency and expedite speed up decision-making, mirroring how society leverages resources for swift outcomes.
VI-B1a Edge-device Collaborative Inference
As depicted in Fig. 25, in edge-device collaborative inference, entities with greater resources offer support to those in need. Different computing layers collaborate to optimize DNN inference processes. The categorization of assistant efforts can be understood through distinct performance optimization objectives:
  • Total Inference Latency Minimization: In EI, minimizing inference latency is analogous to the rapid social response to challenges. Methods like the DINA framework [167] segment DNNs for optimal task allocation, a strategy inspired by social cooperation and matching theory [174]. This emphasis on collaborative problem-solving for reduced response times is also proposed in studies by Chen et al. [175] and Li et al. [176].
  • Inference Accuracy Maximization: Beyond speed, the quality of assistance is vital, reflected in EI’s pursuit of enhanced inference accuracy. Hanyao et al. [177] highlight this with a system that leverages collaborative inference to maintain accuracy under challenging conditions. Similar studies are explored by Yun et al. [178] and Kong et al. [179], emphasizing how shared resources and knowledge can lead to better performance.
  • Total Cost Minimization: Reflecting the social requirement for efficient resource use, EI strives for cost-effective operations. Fang et al. [180] demonstrate this through DNN filter pruning, optimizing resource utilization. This balance is also the focus of research by Du et al. [181] and Dai et al. [182], which mirror social efforts in optimal allocation.
[…]
$\underset{x_{ij}^{d},x_{ij}^{e}}{\text{min}}$ $\sum_{i=1}^{n}\sum_{j=1}^{m}(L_{ij}^{d}+E_{ij}^{d})x_{ij}^{d}+\sum_{i=1}^{k}\sum_{j=1}^{m}(L_{ij}^{e}+E_{ij}^{e})x_{ij}^{e},$ (23)
s.t. $\quad\sum_{i=1}^{n}x_{ij}^{d}+\sum_{i=1}^{k}x_{ij}^{e}=1,\quad\sum_{j=1}^{m}R_{ij}^{d}x_{ij}^{d}\leq R_{di},\quad\forall i,j,$
$\quad\sum_{j=1}^{m}R_{ij}^{e}x_{ij}^{e}\leq R_{ej},\quad x_{ij}^{d},x_{ij}^{e}\in\{0,1\},\quad\forall i,j,$
where $L_{ij}^{d}$ and $L_{ij}^{e}$ are the inference latency of task $t_{j}$ on device $d_{i}$ and edge $e_{i}$, respectively, $E_{ij}^{d}$ and $E_{ij}^{e}$ are the energy consumption of task $t_{j}$ on device $d_{i}$ and edge $e_{i}$, respectively, $x_{ij}^{d}$ and $x_{ij}^{e}$ are the assignment variables indicating whether task $t_{j}$ is assigned to device $d_{i}$ or edge $e_{i}$, $R_{ij}^{d}$ and $R_{ij}^{e}$ are the resource requirements of task $t_{j}$ on device $d_{i}$ and edge $e_{i}$, respectively, and $R_{di}$ and $R_{ej}$ are the available resources on device $d_{i}$ and edge $e_{j}$, respectively.
Edge-device collaborative DNN inference brings about: 1) Lower latency, akin to rapid social assistance. 2) Lower cost, analogous to social efforts aimed at optimizing resource utilization. Despite these advantages, challenges analogous to those in social assistance persist. Limitations in real-time performance reflect social constraints, underscoring that while it introduces a novel paradigm, its evolution is progressive, similar to the evolution of social structures over time.
Figure 26: Illustration of cloud-device collaborative inference.
Communication-efficient Edge-device Collaborative Inference: Edge-device collaborative inference involves frequent communication between ESs and devices, which can be a bottleneck in resource-constrained environments. To address this challenge, various communication-efficient strategies have been proposed, such as model compression, quantization, and selective transmission of intermediate features [183]. These techniques aim to reduce the amount of data exchanged between ESs and devices while maintaining inference accuracy. Additionally, the design of communication-efficient protocols and scheduling algorithms is crucial for optimizing the overall performance of edge-device collaborative inference systems.
VI-B1b Cloud-device Collaborative Inference
Fig. 26 illustrates the collaborative DNN inference within EI, showcasing the dynamic between the cloud and the device that mirrors the social norm of unidirectional assistance. Here, the cloud, rich in resources, empowers devices to enhance their operational efficiency. This interplay is discussed further through the lens of performance optimization objectives:
  • Total Inference Latency Minimization: Vital for real-time applications such as autonomous driving and video surveillance, the balance between accuracy and latency is non-negotiable. Emmons et al. [184] highlight the merit of synergy between cloud and devices. The introduction of complex DNN structures, such as GoogleNet [185] and ResNet [186], complicates the model partitioning. This challenge has been addressed by Hu et al. [187] through the development of an adaptive scheme. Further research by Duan et al. [188] and Jeong et al. [189] emphasizes parallel DNN computation as a solution for latency.
  • Energy Consumption Minimization: While the cloud dominates in computational capacity, device energy limitations are a stark reality. Research by Kang et al. [190] and Hauswald et al. [191] provides strategies for computing partitioning, optimizing both latency and energy use. Collaborative frameworks, as suggested by Eshratifar et al. [192] and Jiang et al. [193], further highlight the need for collaborative frameworks that balance performance with energy consumption.
[…]
where $L=aL_{c}+(1-a)L_{d}$ is the total latency, $E=aE_{c}+(1-a)E_{d}$ is the total energy consumption, $a$ is the partition point of DNN, $L_{c}$ and $L_{d}$ are the latency on the cloud and device, respectively, $E_{c}$ and $E_{d}$ are energy consumption on the cloud and device, respectively, and $\alpha$ and $\beta$ are weights for latency and energy consumption, respectively.
Figure 27: Illustration of cloud-edge-device collaborative inference.
The comparison between assistance and cloud-device collaboration in inference clearly demonstrates similar advantages, such as: 1) Lowered inference latency, achieved through strategic model partitioning and task allocation. 2) Reduced energy consumption, by optimizing the use of communication and computation resources. However, this collaboration also faces distinct challenges: 1) Maintaining performance in unstable conditions, where factors like rapid mobility or communication disruptions can present significant hurdles. 2) Navigating the physical disconnect between cloud and devices, which can limit the scope of applications. This comparison underscores the essential synergy required to achieve superior results in both human society and EI.
VI-B1c Cloud-edge-device Collaborative Inference
Cloud-edge-device collaboration involves specific components (cloud, edge, or device) assisting others without expecting immediate reciprocal benefits. This collaboration branches into three distinct categories, each focusing on enhancing various performance aspects, as shown in Fig. 27.
  • Total Inference Latency Minimization: To address latency issues in collaborative DNN inference systems, Ren et al. [194] formulate a distributed DNN partition. This setup coordinates computations across heterogeneous entities to boost inference efficiency. Further contributions by Xue et al. [195] focus on offloading and distributed DNN deployment, providing beneficial support to the whole system in terms of latency reduction.
  • Total Cost Minimization: Within this collaboration, cost reduction is integral, akin to a form of aid that elevates system functionality. Lin et al. [196] propose an adaptive optimization algorithm, reducing system offloading costs significantly. Similarly, strategies by Teerapittayanon et al. [197] and Liu and Wang et al. [198] develop approaches that strategically map parts of a DNN across the computing hierarchy, optimizing overall costs associated with communication and inference processes.
  • Failure-Resilient Distributed Model: In situations of component failures, the concept of assistance becomes paramount to ensure system robustness. The importance of assistance is portrayed by Yousefpour et al. [199], where skipping hyperconnections in distributed DNNs provides a safeguarding mechanism, ensuring continuity in the face of physical node failures.
[…]
min $\quad L+C-F,$ (25)
s.t. $\quad M=M_{c}\cup M_{e}\cup M_{d},$
$\quad L=\alpha L(M_{c})+\beta L(M_{e})+\gamma L(M_{d}),$
$\quad C=\delta C(M_{c})+\xi C(M_{e})+\pi C(M_{d}),$
$\quad F=\epsilon F(M_{c})+\zeta F(M_{e})+\eta F(M_{d}),$
where $M$ is the entire DNN, $M_{c}$, $M_{e}$, and $M_{d}$ are the partitions of the model executed on the cloud, edge, and device, respectively. $L(M_{x})$, $C(M_{x})$, and $F(M_{x})$ are the latency, cost, and failure resilience of executing partition $M_{x}$ on the corresponding platform, respectively. $\alpha,\beta,\gamma,\delta,\xi,\pi,\epsilon,\zeta,\eta$ are the weights for the respective terms.
Figure 28: Illustration of one-way decision correlation in socialized inference.
Compared to the prior two inference paradigms, cloud-edge-device collaborative inference serves as a supplement to both cloud-device and edge-device collaborative inferences, offering several advantages: 1) Optimized resource utilization, through detailed model segmentation that integrates the full potential of clouds, edges, and devices. 2) Enhanced fault recovery capabilities, by establishing a robust distributed system model that improves sensor fusion, system resilience, and data confidentiality.
Communication Optimization in Cloud-edge-device Collaborative Inference: Cloud-edge-device collaborative inference systems involve coordination and communication among multiple layers, which can introduce significant overhead and complexity. Optimizing the communication efficiency in such systems is crucial for achieving low-latency and high-performance inference. This can be formulated as a multi-objective optimization problem, where the goals are to minimize the communication cost, inference latency, and energy consumption while ensuring the required inference accuracy [200]. Techniques such as hierarchical communication, data compression, and adaptive transmission strategies can be employed to optimize communication efficiency in cloud-edge-device collaborative inference scenarios.
VI-B1d One-way Decision Correlation
The dynamics of one-way decision correlation in EI focus on unidirectional cooperation, as shown in Fig. 28. Decisions made within one hierarchical layer guide subsequent layers, assisting them without direct reciprocation. We explore the implications of unidirectional assistance in one-way decision correlation within EI.
  • Total Inference Latency Minimization: One-way decision correlation streamlines processes by transferring decisions from one layer to another, reducing communication load. Benefiting from upper decisions, lower layers can enhance decision speed and quality, streamlining decision-making and reducing inference latency. Cao et al. [201] present a framework where the cloud ‘big model’ assists edge ‘small models’ by differentiating case complexity, and optimizing task allocation.
  • Energy Consumption Minimization: Similarly, in the energy consumption area, the lower layers align their operations efficiently by receiving and capitalizing on decisions from an upper layer, mitigating unnecessary computational costs. The layer that benefits from this transfer of decisions saves energy, mirroring the dynamic where one party assists another by offering resources or expertise, leading to optimized outcomes. Kosuru et al. [202] showcase this principle within their methods, providing energy-efficient solutions in their respective contexts.
Figure 29: Illustration of edge-cloud co-inference in socialized inference.
[…]
min $\quad\sum_{i}L_{i}(D_{i})+\sum_{i}E_{i}(D_{i}),\ D_{i+1}=f_{i+1}(D_{i}),$ (26)
where $D_{i}$ is the decision made at layer $i$, $f_{i}$ is the decision-making function at layer $i$, $L_{i}(D_{i})$ and $E_{i}(D_{i})$ are the latency and energy consumption of making decision $D_{i}$ at layer $i$, respectively. This non-reciprocal decision transfer aligns with altruism, streamlines processes, reduces inference latency, enhances real-time decision-making for time-sensitive tasks, and conserves energy by optimizing resource use and eliminating unnecessary computations.
VI-B2 Bidirectional Cooperation - Cooperation
Bidirectional cooperation in EI signifies mutual interaction where devices engage in reciprocal information and resource exchange. This interplay ensures a collaborative evolution where individuals learn and adapt based on the broader insights from the EI, leading to more accurate and efficient inferences.
VI-B2a Edge-cloud Co-inference
The edge-cloud co-inference, as shown in Fig. 29, exemplifies bidirectional cooperation in EI. In this scheme, both cloud and edge devices actively participate, leveraging their unique strengths to enhance overall performance and efficiency.
  • Total Inference Latency Minimization: This collaborative dynamic optimizes task allocation based on the inherent capabilities of cloud and edge resources. The cloud, with its superior computing power, handles more complex, resource-intensive tasks, while edges, being closer to the data source, address latency-sensitive tasks. This judicious task allocation offloads certain operations to the cloud, allowing the edge to focus on timely processing, thereby enhancing system adaption. Reinforcing this scheme, Li et al. [203] introduce AppealNet, an architecture aiming to synergize accuracy with low latency in inference. Similarly, Banitalebi Dehkordi et al. [204] advance the collaborative edge-cloud AI through Auto-Split, a mechanism tailoring performance optimization with minimum latency.
  • Energy Consumption Minimization: Energy efficiency is another critical focus of the edge-cloud co-inference. By intelligently offloading tasks to the cloud, edges can operate within their energy constraints, thereby promoting sustainable energy use across the EI system. Further, optimizing resource allocation and power management across this collaborative framework encapsulates the principle of mutual resource conservation. Xiang et al. [205] delve into the cooperation of service provisioning in EC systems, advocating for an energy-efficient scheme. Concurrently, Zhang et al. [206] explore strategies to optimize energy consumption during task offloading in edge-cloud co-inference.
[…]
min $\quad\alpha T_{\text{total}}+\beta E_{\text{total}},$ (27)
s.t. $\quad T_{\text{total}}=T_{\text{edge}}(D)+T_{\text{cloud}}(D)+T_{\text{comm}}(D),$
$\quad E_{\text{total}}=E_{\text{edge}}(D)+E_{\text{cloud}}(D)+E_{\text{comm}}(D),$
$\quad A_{\text{inference}}\geq A_{\text{threshold}},$
where $T_{\text{total}}$ and $E_{\text{total}}$ are the total inference latency and energy consumption, respectively, $T_{\text{edge}}(D)$, $T_{\text{cloud}}(D)$, and $T_{\text{comm}}(D)$ are the latency of executing decisions $D$ on the edge, cloud, and communication, respectively. $E_{\text{edge}}(D)$, $E_{\text{cloud}}(D)$, and $E_{\text{comm}}(D)$ are the energy consumption of executing decisions $D$ on the edge, cloud, and communication, respectively. $A_{\text{inference}}$ is the inference accuracy, $A_{\text{threshold}}$ is the minimum required accuracy. $\alpha$ and $\beta$ are the weights for latency and energy, respectively.
Harnessing both cloud and edge strengths ensures task processing with maximum efficiency, minimum latency and energy conservation. Due to challenges such as network latency, scalability and data privacy, ongoing refinement and adjustment is essential to harness the framework’s potential.
Figure 30: Illustration of edge-device co-inference in socialized inference.
VI-B2b Edge-device Co-inference
Building on edge-cloud collaboration, edge-device co-inference, as depicted in Fig. 30, represents an evolution of bidirectional cooperation in EI, highlighting the synergistic roles of edges and devices. The ensuing discourse delineates the framework of cooperation, underscoring the participatory roles of both edges and devices.
  • Total Inference Latency Minimization: Analogous to social cooperation optimizing task offloading, the edge-device dynamic leverages the strengths of both tiers for latency reduction. Edges, equipped with robust computational resources, absorb significant processing loads, enabling devices to efficiently manage local tasks and circumvent excessive transmission delays. Such an interaction finds support in the work of Bai et al. [207], where the balance between benefits of computational offloading and challenges of communication latency is assessed, highlighting the utility of intelligent reflecting surfaces (IRSs) in enhancing efficiency.
  • Energy Consumption Minimization: Drawing parallels with social emphasis on resource preservation, the edge-device collaboration prioritizes energy conservation. Distributed energy strategies, as seen in Krouka et al. [208], implement model compression techniques for energy optimization at the edge without model performance compromise. Likewise, Xiao et al. [209] propose a dynamic, RL-driven approach, aligning devices to real-time conditions for sustained energy efficiency.
[…]
min $\quad L=L_{e}(T_{e})+L_{d}(T_{d}),\quad E=E_{e}(T_{e})+E_{d}(T_{d}),$ (28)
s.t. $\quad 0\leq T_{e}\leq 1,\quad 0\leq T_{d}\leq 1,\quad T_{e}+T_{d}=1,$
where $L$ and $E$ are the total inference latency and energy consumption, respectively. $L_{e}(T_{e})$ and $L_{d}(T_{d})$ are the latency of executing tasks $T_{e}$ and $T_{d}$ on the edge and device, respectively. $E_{e}(T_{e})$ and $E_{d}(T_{d})$ are the energy consumption of executing tasks $T_{e}$ and $T_{d}$ on the edge and device, respectively. $T_{e}$ and $T_{d}$ are the proportions of tasks executed on the edge and device, respectively. Building on edge-cloud co-inference, edge-device co-inference greater highlights the benefits of bidirectional cooperation. Emphasizing task offloading, adaptive management, and proximity benefits, this mechanism minimizes latency and ensures energy efficiency.
Figure 31: Illustration of decision interaction in socialized inference.
VI-B2c Decision Interaction
Bidirectional decision-making cooperation in EI emphasizes the importance of interdependence between entities, as shown in Fig. 31. This cooperation transcends mere task offloading, focusing instead on collaborative exchange and coordination of decision-making processes and highlighting the interconnectedness of decision-making and the mutual exchange of decision-related insights.
  • Total Inference Latency Minimization: Bidirectional decision-making cooperation aims to minimize inference latency through efficient knowledge pooling. This is achieved by fostering a cooperative environment where devices exchange decision-related insights or intermediate results, synergizing their expertise for timely decisions. An example is presented by Carroll et al. [210], where the FlexiBiT framework illuminates the potential of masked modeling in sequential decision-making. Just as collective decision-making in societies often utilises diverse knowledge sources for better outcomes, this framework’s capacity to be attuned to a range of tasks exemplifies the benefits of collaborative decision-making.
  • Energy Consumption Minimization: In the bidirectional decision-making domain, devices are keenly aware of energy saving. They engage in strategies like intelligent resource allocation or the selective sharing of decision data, all aimed at optimizing energy consumption without compromising decision accuracy. Guo et al. [211] present an exploration into the optimization of energy consumption in distributed inference, consistent with the social principle of strategic resource division for the collective good.
[…]
$\underset{D_{1},D_{2}}{\text{min}}\quad L(D_{1},D_{2})+\lambda E(D_{1},D_{2}),\ \text{s.t.}\quad I(D_{1},D_{2})\geq\theta,$ (29)
where $D_{1}$ and $D_{2}$ are the decision vectors of the two interacting entities, $L(D_{1},D_{2})$ and $E(D_{1},D_{2})$ are the latency and energy consumption of making decisions $D_{1}$ and $D_{2}$, respectively. $I(D_{1},D_{2})$ is the interaction quality between the two decision vectors. $\theta$ is the minimum required interaction quality, and $\lambda$ is the weight for energy consumption.
Decision-making in this cooperative environment focuses on both speed (lower latency) and sustainability (less energy consumption). Emphasizing mutual interaction and coordination underlines the well-known concept that collaboration, whether in society or networks, often exceeds individual efforts. However, to fully benefit from collaborative decision-making, challenges like decision synchronization and information exchange must be managed.
VI-B3 Bidirectional Cooperation - Exchange
In social interactions, exchanging resources and knowledge is crucial for mutual benefit. This principle, when applied to EI, entails an interaction between hierarchical layers during the inference phase. The cloud, with its extensive data repositories, requires data from edges for context-aware inferences, while edge layers use the cloud’s computing power for deeper analysis. This interplay ensures inferences are prompt and relevant, and capitalize on each stratum’s strengths, fostering an integrated and efficient system. We explore how clouds, edges, and devices while maintaining their distinct objectives, engage in this reciprocal exchange during task inference.
  • Total Inference Latency Minimization: Reciprocity in bidirectional exchanges aids in optimizing inference latency. While the cloud focuses on depth, the edge emphasizes speed and context. Devices, being closest to the source of data, prioritize immediate, low-latency inferences, drawing insights from the edge or cloud. The aim is a harmonious distribution of tasks that benefits from each layer’s strengths. Li et al. [212] delve into reducing latency, particularly during simultaneous inference across multiple models with resource constraints. The research pivots to task scheduling in real-time, utilizing the ONNX runtime engine for inference. It highlights the challenges in lowering system latency and maintaining compatibility across different platforms, particularly during simultaneous model execution.
    TABLE XI: Comparative Analysis of Inference Approaches Reflecting Cooperation Interactions
    Type
    Social
    Relationship
    Technique Efficiency Scalability
    Communication
    Overhead
    Heterogeneity
    Handling
    Cost
    Efficiency
    Intra - layer Mutual Assistance
    Edge-edge
    Collaborative
    Inference
    [166, 167, 168]
    Optimized for
    minimum latency
    Highly scalable
    due to
    decentralization
    Moderate due to
    inter-edge
    communication
    Challenges due
    to diverse edge
    Costs saved by
    reduce server
    but increase with
    communication
    overhead
    Device-device
    Collaborative
    Inference
    [170, 171, 172, 173]
    Balances between
    transmission and
    computational
    latency
    Scalable but
    dependent on
    device
    capabilities
    High due to
    frequent
    communication
    Sync delays
    due to
    varied device
    Costs saved on
    centralized
    processing
    Inter - layer Assistance
    Edge-device
    Collaborative
    Inference [167],
    [174, 175, 176, 177, 178, 179, 180, 181, 182]
    Optimized by
    task allocation
    High with edge
    scalability
    Moderate due to
    edge-device
    communication
    High due to
    diverse device
    High with
    optimized
    resource use
    Cloud-device
    Collaborative
    Inference
    [184, 185, 186, 187, 188, 189, 190, 191, 192, 193]
    Optimized by cloud’s
    computing power
    High with cloud
    scalability
    High due to
    cloud-device
    distance
    Moderate
    Moderate due
    to cloud costs
    Cloud-edge
    -device
    Collaborative
    Inference
    [194, 195, 196, 197, 198, 199]
    High with multi
    -layer
    optimization
    Very High
    High with
    multi-layer
    communication
    Very High
    High with
    optimized
    resource
    distribution
    One-way
    Decision
    Correlation
    [201, 202]
    Streamlined by
    decision transfer
    Moderate
    Low due to
    one-way transfer
    Moderate
    High with
    reduced
    redundancy
    Cooperation
    Edge-cloud
    Co-inference
    [203, 204, 205, 206]
    High with shared
    responsibilities
    Very High
    Moderate due to
    edge-cloud proximity
    High
    High with
    resource sharing
    Edge-device
    Co-inference
    [207, 208, 209]
    High with
    localized
    processing
    High with
    edge scalability
    Moderate due to
    edge-device
    proximity
    Very High
    High with
    localized
    processing
    Decision
    Interaction
    [210, 211]
    High with mutual
    decision insights
    High
    High due to
    continuous
    interaction
    High
    Moderate with
    decision overheads
    Exchange -
    High with device-
    level optimization
    Moderate with
    device limits
    Moderate due to
    device-device
    proximity
    High with
    device
    compatibility
    High with
    device-
    level optimization
  • Energy Consumption Minimization: The principle of bidirectional exchange also applies to energy consumption. Each layer aims to use energy efficiently in making inferences. The cloud leverages its vast resources, offloading tasks to the edge to conserve energy. The edge, due to its proximity to data sources, minimizes energy consumption during data transmission and focuses on local inferences. Devices, inherently resource-constrained, advocate for energy-efficient inferences and selective data transmissions. Hu et al. [213] spotlight energy conservation in mobile networks, illustrating how edge situated closer to mobile devices can optimize energy consumption. When a device’s computational resources are insufficient, offloading tasks to the edge presents an energy-conservative solution.
[…]
max $\quad\theta_{C}O_{C}(X_{C})+\theta_{E}O_{E}(X_{E})+\theta_{D}O_{D}(X_{D})$ (30)
$-\lambda[L_{c}(X_{C})+L_{e}(X_{E})+L_{d}(X_{D})]$
$-\omega[E_{c}(X_{C})+E_{e}(X_{E})+E_{d}(X_{D})],$
s.t. $\quad X_{C}+X_{E}+X_{D}\leq T.$
Task and data exchanges between the cloud ($X_{C}$), edge ($X_{E}$), and devices ($X_{D}$) are optimized in our model to meet the objectives of each ($O_{C}$, $O_{E}$, $O_{D}$), while considering latency ($L_{c}$, $L_{e}$, $L_{d}$) and energy consumption ($E_{c}$, $E_{e}$, $E_{d}$). Weights ($\theta_{C}$, $\theta_{E}$, $\theta_{D}$, $\lambda$, $\omega$) prioritize these factors accordingly, ensuring that the total exchange ($T$) aligns with system capabilities and objectives. This inter-layer collaboration, mirroring social reciprocity, improves task allocation efficiency, reduces latency and conserves energy by leveraging participant’s strengths.
The discussion on Socialized Inference in EI draws an analogy between social cooperation and collaborative decision-making, such as edge-device and cloud-edge interactions. Detailed in Table XI, these mechanisms manage the trade-off between latency and energy efficiency, some focusing on real-time response and others on energy savings.
VI-C Intra-layer Submissive Interaction in Socialized Inference
VI-C1 Conformity
Similar to how individuals in society align their actions with group norms to promote cohesion, devices also adhere to established norms to enhance efficiency.
VI-C1a Ultra-Reliable and Low-Latency Communication
Conformity in EI focuses on enabling ultra-reliable and low-latency communication (URLLC) by streamlining processes and eliminating superfluous exchanges to facilitate swift and reliable decision-making. To encapsulate these ideas into a formula, we can represent the system’s efficiency as a function of conformity, latency, and reliability. The SL function Eq. (1) for URLLC is $F([\text{devices}],[\text{conformity}],\varnothing,\text{inference},\text{tasks})$. The optimization objective is:
max $\quad U(C,L,R)=C\times R-\alpha\times L,$ (31)
s.t. $\quad 0\leq C\leq 1,\quad L\geq 0,\quad 0\leq R\leq 1,$
where $C$ is the conformity among devices, $L$ is the communication latency, $R$ is the reliability, $U(C,L,R)$ is the utility function. Constant $\alpha$ weights the significance of latency, balancing it against reliability.
Han et al. [214] present “coded edge federation” (CEF), where edge infrastructure providers (EIPs) join forces to allocate resources for coded distributed computing (CDC) tasks. Their approach shows how conformity leads to a Nash equilibrium in resource allocation, with no benefit from unilateral strategy changes. Tran et al. [215] highlight how conformity to protocols in 5G networks enhances EI system performance.
Figure 32: Illustration of task-oriented communication in socialized inference.
VI-C1b Task-oriented Communication
In human societies, conformity impacts cooperative task execution. Likewise, in EI, conformity dynamics serve as the backbone for task-oriented communication. Devices streamline communications to be task-centric. Conformity ensures interactions align with collective task completion goals. The shared understanding that arises from conformity forms the foundation for devices to synchronize their activities and share task-relevant information, as shown in Fig. 32. It facilitates collaborative decision-making and reduces potential for communication conflicts. The SL function Eq. (1) for task-oriented communication is $F([\text{devices}],[\text{conformity}],\varnothing,\text{inference},\text{communications})$. The optimization objective is:
${\text{max}}\quad C\times f-\alpha\times L,\quad\text{s.t.}\quad 0\leq C,f\leq 1,L\geq 0,$ (32)
where $C$ is the conformity among entities, $f$ is the relevance of features to the task, $L$ is the communication latency, and $\alpha$ is the weight for latency.
Shao et al. [216] delve into this paradigm by investigating task-oriented communication in multi-device cooperative edge inference. Distributed edge devices transmit local sample features to a more powerful edge for inference. While it enhances the collective sensing capability beyond that of a single device, it does introduce challenges, notably increased communication overheads and potential latency. However, Shao et al. navigate this challenge by proposing a learning-based communication model that optimally extracts local features and encodes distributed features in a strictly task-oriented manner.
VI-D Inter-layer Submissive Interaction in Socialized Inference
VI-D1 Unidirectional Submissive Interaction - Suggestion&imitation
In traditional societies, authoritative figures guide behaviors. The suggestion&imitation pattern, depicted in Fig. 33, shows how leaders guide actions. In EI inference, layers with contextual or computational advantages provide ‘suggestions’ as processed insights. Other layers then ‘imitate’ by aligning their strategies with these insights. Unlike model parameter and feature transfer in Socialized Training, Socialized Inference focuses on real-time decision alignment, harmonizing actions with guiding layers. This ensures EI cohesion, reflecting how authoritative figures influence group decisions. We show how different layers’ dynamics affect communication latency, energy usage, and computation.
Figure 33: Illustration of suggestion & imitation in socialized inference.
  • Resource Allocation: Resource allocation in EI, emulating social dynamics, involves the harnessing of resources and the imitation of efficient strategies. Authoritative layers guide resource allocation by suggesting optimal routes or efficient edge strategies, steering lower layers towards optimization and reducing communication delays. Cen et al. [217] explore an RL approach to these challenges in cloud-edge environments. Similarly, Hu et al. [218] investigate intelligent resource allocation in edge-cloud networks using a hybrid DDPG-D3QN model.
  • Energy Management: Authoritative layers’ suggestions drive energy-efficient strategies, urging edges to mimic energy-saving techniques or adopt operational modes. Such imitation dynamics prolong device lifespans, reducing overall energy wastage. Ruan et al. [219] propose a three-tier architecture, emphasizing latency improvements and energy-efficient scheduling.
  • Computational Offloading: Task balancing between cloud and edge is central to EI. Authoritative entities steer lower layers’ offloading decisions for balanced computational loads, reducing communication latency and enhancing system responsiveness. Wang et al. [220] focus on computational offloading in wireless networks, balancing ES capacity and energy consumption. Zeng et al. [221] present a joint optimization for computational offloading, assessing its effects on energy usage and performance.
[…]
max $\quad\theta_{1}R(S,I)+\theta_{2}E(S,I)+\theta_{3}O(S,I),$ (33)
s.t. $\quad 0\leq S,I\leq 1,$
where $S$ and $I$ are the levels of suggestion and imitation, respectively, $R(S,I)$, $E(S,I)$, and $O(S,I)$ are the performance metrics for resource allocation, energy management, and computational offloading, respectively, and $\theta_{1}$, $\theta_{2}$, and $\theta_{3}$ are the weights for the respective metrics.
TABLE XII: Comparative Analysis of Inference Approaches Reflecting Submissive Interactions
Type
Social
Relationship
Technique Efficiency Scalability
Communication
Overhead
Heterogeneity
Handling
Cost
Efficiency
Intra - layer Conformity URLLC [214, 215]
Optimized due to
streamlined
communication
High due to
uniformity
Reduced due to
streamlined
channels
Moderate due to
shared standards
Optimized due to
reduced redundancies
Task-oriented
Communication
[216]
Enhanced due to
task-centric focus
High due to
shared task
objectives
Increased due to
task-specific
exchanges
High due to
task-specific
handling
Optimized due to
task-centric
focus
Inter
layer
Suggestion
Imitation
[217, 218, 219, 220, 221]
Optimized by
task allocation
Enhanced due to
authoritative
suggestions
Reduced due to
streamlined
suggestions
High due to
centralized
guidance
Optimized due to
authoritative
suggestions
Communication Efficiency in Suggestion&Imitation Mechanisms: Communication efficiency plays a crucial role in ensuring effective information transmission from higher to lower layers in Suggestion&Imitation mechanisms. To optimize end-to-end communication performance, various strategies can be employed, such as information compression, quantization, and selective transmission. These techniques aim to reduce data transferred between layers while maintaining inference accuracy. Moreover, designing efficient communication protocols and scheduling algorithms is essential for minimizing latency in the suggestion and imitation processes [222]. By optimizing communication efficiency in Suggestion&Imitation mechanisms, socialized inference systems can achieve faster convergence and lower resource consumption.
Unidirectional submissive interactions, like suggestion and imitation dynamics, are evident in both human societies and EI inference, emphasizing the potential benefits of such collaborative frameworks. Embracing these dynamics enables a harmonized, efficient, and responsive system, optimizing resource allocation, energy management, and computational tasks. This hierarchical structure optimizes processes from resource allocation to energy management. A comparative analysis presented in Table XII elucidates interactions’ distinctions and synergies.
Figure 34: Illustration of competition in socialized inference.
VI-E Intra-layer Conflict Interaction in Socialized Inference
VI-E1 Competition
In societies, competition can drive innovation but may also cause conflicts. This duality is mirrored in EI’s inference, where devices within the same layer vie for superiority as illustrated in Fig. 34. Competitors in inference rely more on quick decisions and resource optimization than dataset dominance. The key lies in efficient responses without compromising EI’s core objectives. Managing intra-layer competition effectively becomes essential to strike a balance between performance improvements and optimizing system resources and capabilities.
  • Resource Allocation: Intra-layer competition centers on resource allocation. Devices competing for limited computational resources may increase communication demands, adversely affecting communication latency. Additionally, increased competition often leads to higher energy demands, compromising the energy efficiency of the system. Fan et al. [168] investigate networks where agents compete for resources, proposing a mechanism for efficient and fair allocation. Sun et al. [223] optimize for maximum network satisfaction through fair resource sharing.
  • Task Offloading: In edge-cloud collaborations, task offloading becomes another aspect of intra-layer competition. Devices jostle for offloading their tasks to available, capable nodes, potentially affecting latency and energy consumption. Offloading strategies can either promote energy efficiency or, if unchecked, exacerbate energy demands. Ji et al. [224] contribute to this issue by focusing on energy-efficient offloading algorithms, considering latency tolerances and uncertainties inherent in such networks.
  • Computing Efficiency: The competition is also driving devices to bolster their computing efficiency. Devices vie to optimize algorithms, employ cutting-EC methods, or adopt strategies that promise efficiency. Such pursuits, while fostering innovation, could inadvertently increase computational overheads. Lim et al. [225] provide insight into scheduling algorithms in allocation scenarios and offer effective strategies for task execution in competitive environments. Zhang et al. [226] expand on this by discussing computing schedules in EI that are energy-efficient.
[…]
max $\quad E=\lambda_{1}R(C)+\lambda_{2}T(C)+\lambda_{3}Comp(C),$ (34)
s.t. $\quad 0\leq C\leq 1,\quad 0\leq R(C),T(C),Comp(C)\leq 1,$
$\quad\lambda_{1}+\lambda_{2}+\lambda_{3}=1,$
where $C$ is the level of competition, $R(C)$, $T(C)$, and $Comp(C)$ are the efficiency metrics for resource allocation, task offloading, and computation, respectively, and $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are the weights for the respective metrics.
The concept of competition is not only central to society but also critical within the domain of inference in EI. Devices within the same layer strive to excel in resource allocation, task offloading, and computing efficiency while maintaining a balance with the collective energy and computational needs of the system. The above works emphasize how important it is to manage this competition and offer methods for coordinating personal objectives with the overall objectives of EI in order to ensure reliable and effective conclusions.
Communication-aware Resource Allocation in Competitive Inference: In competitive inference scenarios, where multiple devices or agents compete for shared resources, communication efficiency plays a critical role in determining the overall system performance. The communication network topology, bandwidth allocation, and transmission strategies need to be carefully designed and optimized to ensure fair and efficient resource allocation among competing devices. Game-theoretic approaches and distributed optimization techniques [227] can be applied to model and solve communication-aware resource allocation problems in competitive inference settings.
Figure 35: Illustration of compulsory in socialized inference.
VI-F Inter-layer Conflict Interaction in Socialized Inference
VI-F1 Unidirectional Competition - Compulsory
In social hierarchies, upper echelons dictate the actions of subordinate layers, establishing a unidirectional command chain, as depicted in Fig. 35. Although this hierarchy centralizes decision-making and optimizes inference, it also generates issues such as communication delays, energy demands, and computing inefficiency. Mitigating these challenges necessitates collaborative strategies in resource allocation, task offloading, and scheduling, thereby enhancing resource efficiency, communication efficiency, and overall performance.
  • Resource Allocation: The advantage in resource allocation is evident when higher strata enforce allocation decisions upon inferior layers. While it offers centralized control and enhanced resource optimization across the edge-cloud spectrum, it also brings about concerns related to communication latency, instigated by coordination overheads. Collaborative resource allocation algorithms, as studied by Jiang et al. [228], can act as mitigators to combat this. Furthermore, energy conservation gives rise to the development of energy-centered allocation strategies that target efficient resource allocation and energy consumption.
  • Task Offloading: In the edge-cloud collaboration, unidirectional compulsory is manifest when superior layers necessitate task offloading to their lower counterparts. This mechanism enhances centralized management and scalability but also raises latency issues. Almutairi et al. [66] delve into task offloading, highlighting a fuzzy logic system to optimize the end-to-end service time. Proximity-based communication and energy-efficient offloading ensure tasks find their way to energy-optimized nodes.
  • Computing Efficiency: Unidirectional compulsory within layers during inference accentuates the edge-cloud collaborative system’s computing efficiency. Although this compulsory can be demanding, it also catalyzes advancements in performance. Given the heterogeneity of computational resources, collaborative algorithms and ML approaches are needed. Maray et al. [229] introduce a dynamic time-sensitive scheduling algorithm, while Fan et al. [230] aim at optimizing the total task processing delay. Furthermore, Zhang et al. [231] bring forth a mechanism optimizing the task completion time, emphasizing the limited storage space of the ES and its implications.
[…]
max $\quad\delta_{1}\cdot R(H)+\delta_{2}\cdot T(H)+\delta_{3}\cdot Comp(H),$ (35)
s.t. $\quad 0\leq H\leq 1,$
where $H$ is the level of directive from higher-level servers, $R(H)$, $T(H)$, and $Comp(H)$ are the efficiency metrics for resource allocation, task offloading, and computation under the higher-level directive, respectively, and $\delta_{1}$, $\delta_{2}$, and $\delta_{3}$ are the weights for the respective metrics.
TABLE XIII: Comparative Analysis of Inference Approaches Reflecting Conflict Interactions
Type
Social
Relationship
Efficiency Scalability
Communication
Overhead
Heterogeneity
Handling
Cost
Efficiency
Intra
layer
Competition
[168, 223, 224, 225, 226]
Optimized due to
intra-layer rivalry
fostering innovation
Moderate due to
potential resource
contention
Elevated due to
frequent resource-
negotiation
exchanges
Diverse due to
individual entities’
unique strategies
Balanced due to
competitive
resource utilization
Inter - layer
Compulsory
[228, 66, 229, 230, 231]
Centralized due to
top-down directives
optimizing processes
High due to
structured and
directed resource
allocation
Streamlined due to
clear directives
minimizing
back-and-forth
Uniform due to
centralized standards
and guidelines
Economical due to
optimized and
directed resource
allocation
Struggle
[232, 233, 234, 235, 236, 237]
Adaptive due to
mutual adjustments
and negotiations
Dynamic due to
continuous
inter-layer
adjustments
Intensive due to
constant
bidirectional
negotiations
Varied due to
mutual influence
and diverse strategies
Negotiable due to shared
responsibilities
and mutual
adjustments
Figure 36: Illustration of struggle in socialized inference.
VI-F2 Bidirectional Competition - Struggle
In society, struggles involve complex interactions as people pursue objectives while reacting to others’ actions (as shown in Fig. 36). In inference, layers face bidirectional struggles over shared resources or conflicting goals. Similar to how social struggles inspire innovation, these interactions might lead to optimal solutions. However, unmanaged conflicts run the risk of inefficiencies.
  • Resource Allocation: Within EI, resources are finite, and the allocation dynamics mirror the social struggle for scarce resources. Each layer competes for these resources, driving a push towards greater operational efficiency. However, this struggle can also result in increased latency and energy consumption. Collaborative communication protocols, such as those proposed by Lee et al. [232], aim to enhance transparent communication and improve efficiency. Concurrently, energy consumption concerns are tackled by methods like those suggested by KN et al. [233], which seek to balance efficiency with assured QoS.
  • Task Offloading: The social dynamics of responsibility-sharing and allocation is reflected in the act of task offloading. Layers struggle with decisions regarding task offloading, often influenced by their intrinsic capabilities and external pressures. The challenge lies in balancing between local processing and offloading, each with its pros and cons. Nguyen et al. [234] offer an illustrative approach by promoting UAV collaboration, to achieve optimal offloading. Similarly, strategies presented by Wang et al. [235] promote a multi-tiered approach, echoing the social hierarchies and their interactions.
  • Computing Efficiency: Efficiency in computations, much like productivity in society, can be threatened by internal frictions or conflicts. Struggle interactions can give rise to redundant computations or misaligned task allocation. To address these issues, systems can emulate social cooperation frameworks. Algorithms, as proposed by Xavier et al. [236], mimic social cooperation to ensure optimal resource utilization. Amer et al. [237] introduce a hierarchical approach to task scheduling, mirroring social hierarchies, with each level having its own set of responsibilities and interactions.
[…]
max $\quad\eta_{1}R(B)+\eta_{2}T(B)+\eta_{3}Comp(B),$ (36)
s.t. $\quad 0\leq B\leq 1,$
where $B$ is the level of bidirectional struggle, $R(B)$, $T(B)$, and $Comp(B)$ are the efficiency metrics for resource allocation, task offloading, and computation under bidirectional struggle, respectively. $\eta_{1}$, $\eta_{2}$, and $\eta_{3}$ are the weights for the respective metrics.
Communication Resource Competition in Struggle Interactions: In Struggle interactions, devices or agents from different layers compete for shared communication resources, which can lead to degraded inference performance. To mitigate this issue, effective communication resource allocation strategies need to be designed. One possible approach is to utilize game theory and optimization techniques to model and solve the communication resource allocation problem [238].
The objective is to fairly and efficiently allocate communication bandwidth and time slots among struggle devices, considering their inference task requirements and priorities. Additionally, hierarchical communication resource management frameworks can be employed to coordinate resource allocation across different layers, reducing conflicts and improving overall inference performance. By optimizing communication resource allocation in struggle interactions, socialized inference systems can achieve more stable and efficient inference.
The bidirectional struggle in EI reflects the complex interdependencies familiar in society, emphasizing the continuous pursuit of resources, task offloading efficiency, and computing power. Efficiency and complexity are stimulated by the struggle that exists within EI. EI’s struggle, which is marked by rivalries at many levels, encourages creativity but makes energy management and communication more challenging. Table XIII describes conflict interactions in-depth and emphasizes how they affect the scalability and efficiency of the system.
Figure 37: The taxonomy graph of the unified SL function. It maps the diverse methods of socialized training and inference in EI to the variables $(\mathbf{A},\mathbf{R}_{w},\mathbf{R}_{a},\mathbf{I},\mathbf{L})$ defined in Eq. (1). This enables researchers to design customized and optimized SL schemes by selecting appropriate values for these variables, tailored to specific EI application requirements.
VI-G Lessons Learned
Lessons learned section highlights how social principles can improve decision-making and the overall adaptability of EI. EI is moving from isolated operations to increased collaboration, mirroring cooperative decision-making observed in social dynamics. Similar to social governance trends, EI is moving toward decentralized decision-making. Similar to social resource management, the latency-energy trade-off balances energy savings with timely EI responses. Additionally, EI’s collaboration-competition dichotomy influences how resources are allocated and how quickly decisions are made, reflecting social dynamics.
To ensure seamless interactions within EI, protocols and standardization are necessary, as they align with social norms and regulations. Similar to social issues, protecting sensitive information requires a high priority for security and privacy as EI becomes more interconnected. The efficiency of EI depends on advancing algorithms and technology to address these issues. This technological evolution in EI parallels social advancements, suggesting a future where EI’s complexity is adeptly aligned with the multifaceted nature of society.
In summary, the taxonomy graph of the SL function, presented in Fig. 37, encapsulates the various aspects of socialized training and socialized inference discussed throughout Sections V and VI, each mapped to the SL function defined in Eq. (1). The graph illustrates the diverse range of SL customization and optimization possibilities in EI scenarios. By carefully selecting the appropriate values for these variables, researchers and practitioners can design and implement SL schemes tailored to their specific EI application requirements. The taxonomy graph serves as a comprehensive reference for understanding the scope and potential of SL in EI systems, guiding future research and development efforts.
VII Tutorials for Implementing SL in EI
To effectively realize SL functions and ensure their seamless integration into EI systems, the development and adoption of standardizations and platforms are crucial. Standardizations provide a common framework for implementing SL in EI, addressing critical aspects such as data formats, communication protocols, and security mechanisms. Platforms, on the other hand, offer the necessary software tools, libraries, and hardware components to support the deployment and execution of SL algorithms in EI environments. In this section, we discuss the key standardizations and platforms that enable the efficient implementation of SL in EI systems.
VII-A Standardizations for Implementing SL in EI
VII-A1 Learning
The IEEE’s FL standard IEEE 3652.1-2020 [239], developed by industry leaders, sets a foundational architecture for SL in EI. The importance of this standard lies in its ability to provide a common framework for secure and scalable FL, which is a key enabler for SL in EI. By defining guidelines for data governance, model management and system architecture, IEEE 3652.1-2020 facilitates deployment of FL in various EI applications, such as smart healthcare and industrial IoT. It addresses critical aspects like data, user, and system requirements, enabling secure and scalable SL-driven applications. Further, general standards for privacy and security in IoT and smart cities, like ISO/IEC TS 27570 [240] and ISO/IEC DIS 27400 [241], are important. These standards, along with consumer IoT cybersecurity baseline European Telecommunications Standards Institute (ETSI) EN 303 645 [242], reinforce framework for trustworthy SL solutions in EI.
VII-A2 Computing
In modern wireless networks, the shift towards digital modulation has propelled MEC as a vital component in deploying EI systems [47]. The ETSI ISG MEC aims to standardize a system conducive to EC, laying the foundation for SL in EI [243]. The 3GPP 5G specifications lay the groundwork for EC integration [244], and the advancement toward 6G opens prospects for unifying communication and computing, potentially integrating AirComp functionalities for enhanced edge training [245]. SL’s adaptive and responsive nature will be crucial in these evolving systems, ensuring their alignment with social trends and user interactions.
VII-B Platforms for Implementing SL in EI
VII-B1 Software
The advancement of EI has pushed the development of a variety of software platforms to simulate and operationalize EI algorithms. Frameworks like TensorFlow Federated and PySyft provide the foundational groundwork for EI simulations. Platforms such as FedML [246], FATE [247], KubeEdge, and HarmonyOS [248] extend these capabilities, offering versatile libraries and secure computing architectures for a broad range of EI applications, from retail to healthcare. For instance, the FedML platform offers a comprehensive framework for FL, supporting a wide range of algorithms, models, and datasets. It enables researchers and practitioners to easily develop and deploy FL solutions in EI environments, accelerating the adoption of SL in real-world applications. The platform’s modular architecture and extensive documentation make it accessible to users with varying levels of expertise, fostering collaboration and innovation in the field of SL and EI. These platforms play a crucial role in bridging the research-production gap and enhancing real-time EI applications, setting the stage for the integration of SL, which further refines EI through continual user interaction and social inputs.
Figure 38: Future applications of SL in EI.
VII-B2 Solver
In the expanding landscape of EI and wireless networks, resource allocation optimization tools become increasingly critical. Tools like CVX [249], MOSEK [250], and Gurobi [251] are pivotal for large-scale convex optimization and handling mixed-integer conic programs. Techniques like the matrix stuffing technique [252] accelerate the modeling process, enhancing the efficiency of symbolic conic transformation solvers. The integration of these solvers with SL paradigms on EI platforms is essential. An example is the “Open-L2O” framework [253], which blends learning to optimize algorithms, showcasing how SL can be instrumental in adapting EI based on user interactions and feedback.
VII-B3 Hardware
EI’s performance hinges on the interplay between specialized computing hardware like NVIDIA’s GPUs and Google’s TPUs and RF technologies [47]. Advances such as deep RL-assisted chip floorplanning [254] and IRS-enhanced FL systems [255] highlight the evolution in EI hardware design. Additionally, emerging THz communication [256] is set to unify sensing, communication, and learning, further propelling EI capabilities. Incorporating SL is essential to ensuring that EI remains adaptable and user-centric, continually improving its efficiency and responsiveness.
VIII Future Applications and Future Research
VIII-A Future Applications of SL in EI
In Section I-C, we provided a concise overview of the current application scenarios of SL in EI, highlighting the common characteristics and potential benefits across various domains. This section provides an in-depth analysis of SL in EI’s future applications, exploring emerging domains and technologies. We comparatively study these application scenarios, discussing their unique requirements, challenges, and opportunities. As illustrated in Fig. 38, the integration of SL into EI has an impact on multiple future application scenarios. This section examines various uses and offers guidance on how to put them into practice.
Digital Twins. Digital twins, as digital representations of physical entities, are evolving to incorporate SL for enhanced adaptability and optimization. Implementing complex digital twin models efficiently on resource-constrained edge devices is a key challenge. Future research should focus on developing lightweight, yet high-precision physical models and distributed training techniques. For instance, Lu et al.’s DTWN [257] could be extended to incorporate FL, enabling collaborative model updates across multiple edge nodes while preserving data privacy. Sun et al.’s task offloading approach [258] could be enhanced with RL-based decision-making to dynamically optimize offloading strategies based on network conditions and task characteristics. Future work in space-air networks [259] could explore hierarchical digital twins leveraging SL for multi-level optimization from individual device management to global network orchestration. Bellavista et al.’s resource allocation research [260] could be expanded to include predictive analytics, utilizing SL to anticipate and proactively adjust to resource demands.
Web 3.0. The integration of EI and Web 3.0 promotes the creation of decentralized applications (DApps) on IoT edges [261], which poses unique challenges for SL. Developing efficient consensus mechanisms in decentralized environments is a critical issue. Future research should investigate SL-enhanced consensus algorithms, like reputation-based systems where node credibility hinges on SL performance. Integrating SL in DApp development could enable adaptive smart contracts that evolve based on user interactions and network conditions. This necessitates novel contract languages that handle probabilistic and learning-based clauses. Privacy-preserving SL techniques require optimization on resource-constrained edge devices to ensure data security in EI environments [262]. Research in industrial automation should focus on real-time SL algorithms for machine-to-machine interactions, employing online learning techniques to adapt to changing production conditions while upholding the robustness of Web 3.0.
Computing Power Network (CPN). Integrating EI into CPN presents optimization challenges for achieving ultra-low latency and high bandwidth, especially near base stations [263]. Future studies should develop dynamic resource allocation algorithms that adjust in real time to SL demands. This approach requires predictive resource allocation techniques that utilize learned SL task patterns. Designing distributed learning algorithms that efficiently utilize heterogeneous computing resources across the CPN is a key challenge. As industries demand customization [264], research must explore adaptive SL algorithms that dynamically adjust to varying computing resources and network conditions. The shift towards 6G micro-edge architectures demands novel SL frameworks that operate seamlessly across diverse EI sites, leveraging FL and others to balance data locality with global model optimization.
Blockchain. Although the integration of Blockchain and EI optimizes IoT systems [49, 265], it presents challenges in efficiency and scalability. Future studies should develop SL-enhanced consensus mechanisms tailored for resource-constrained edge devices. Liang et al.’s work on obstacle detection in train systems [266] could adopt FL to enable collaborative model updates while preserving data privacy across distributed nodes. Designing SL algorithms that leverage blockchain’s immutability for secure, decentralized learning presents a key challenge. Building on Sharma et al.’s IoT security framework [267], future research should explore using distributed SL in blockchain-based anomaly detection to predict threats from historical node interactions. To enhance scalability in Desai et al.’s system [268], research should focus on combining sharding techniques with SL to boost throughput in large-scale IoT networks.
AIGC. Integrating SL into AIGC enhances its ability to adaptively learn from user interactions and collective insights, personalizing content, but also introduces unique challenges beyond traditional approaches [269]. Future research should focus on SL-enhanced generative models that adapt to collective user behaviors at the edge. Building on Xu et al.’s multi-layered framework [270], developing SL algorithms that facilitate collaborative learning of GANs and VAEs across distributed edge nodes is necessary. Designing privacy-preserving SL techniques for personalized content generation, which potentially leverage FL with differential privacy, is a critical challenge. Research should focus on exploring SL-driven adaptive model partitioning that dynamically allocates tasks between different layers, based on social patterns and device capabilities. Furthermore, investigating SL-based content optimization techniques at the edge for various modalities, such as text [271], images [272], and videos [273], while balancing quality, latency and resource constraints, is crucial.
Quantum Information. Integrating quantum computing principles with SL in EI presents unique challenges for enhancing distributed learning and security. Research should prioritize developing quantum-inspired SL algorithms to enhance collaborative learning efficiency in edge networks. For path planning [274], quantum-inspired optimization techniques could enable edge devices to collaboratively solve complex problems more efficiently. Designing privacy-preserving SL protocols that leverage quantum-resistant cryptography to ensure long-term security against advancing quantum capabilities remains a key challenge. Investigating quantum key distribution (QKD) for securing SL data exchanges in edge networks is crucial [275]. Future work should aim to integrate quantum-safe cryptography within SL frameworks, focusing on scalability and resilience in real-world edge environments.
Metaverse. The Metaverse, which merges virtual and real worlds through 5G, AI, and VR technologies, presents unique challenges for EI and SL integration [276]. As Metaverse complexity grows, the trend towards EC becomes crucial for improving performance and reducing latency. Extending Younis et al.’s edge-cloud architecture [277], future research should focus on SL-enhanced bandwidth optimization and data security in Metaverse environments. Key challenges include developing adaptive SL algorithms for dynamic edge-cloud resource allocation based on real-time user interactions and network conditions. Research into SL techniques that efficiently synchronize heterogeneous data types (visual, auditory, haptic) across distributed edge nodes in the Metaverse is needed. Furthermore, exploring SL-driven content adaptation mechanisms is crucial for developing virtual environments that evolve in response to collective user behaviors.
AR and VR. AR/VR require low-latency interactions, which are facilitated by edge-cloud technologies [278]. Future studies should develop SL-enhanced task allocation algorithms for hybrid cloud-edge-device architectures, focusing on Mobile AR [279]. Designing adaptive SL models that dynamically allocate computational tasks between edge and cloud, responding to real-time user interactions and network conditions, is a major challenge. Given intense data demands and latency sensitivity, investigating SL techniques for predictive content caching and rendering at the edge is crucial [280]. Research into collaborative SL algorithms that facilitate real-time content adaptation across multiple AR/VR devices, leveraging collective user interactions to create personalized experiences, is needed. Furthermore, exploring SL-driven compression techniques for data streams could reduce bandwidth requirements while preserving QoE.
Comparative Analysis of SL in EI Application Scenarios. Integrating SL in EI across various domains highlights both commonalities and unique challenges. While collaborative learning, adaptability, privacy preservation and others are core principles of SL in EI, their implementation varies significantly depending on application-specific requirements. Data characteristics and real-time processing demands vary significantly across domains: digital twins and metaverse applications process complex, multi-modal data streams, whereas Web 3.0 and blockchain systems handle distributed, often encrypted data. This variability necessitates developing adaptive SL frameworks that can efficiently handle heterogeneous data types and processing requirements.
Device heterogeneity presents another critical challenge, as the computational capabilities range from ESs in CPNs to resource-constrained IoT devices in AIGC scenarios. Future research should develop scalable SL algorithms capable of dynamically adjusting to diverse device capabilities, utilizing techniques like model compression, adaptive quantization, and intelligent task offloading. The objectives of SL also differ across applications: AR/VR systems prioritize low-latency content adaptation, whereas quantum-enhanced EI focuses on secure, distributed computation.
Despite these differences, it is crucial to maintain SL’s core principles across all applications. Future research should investigate unified SL frameworks that can adapt to diverse scenarios while ensuring collaborative learning, system adaptability, and privacy preservation. This approach could involve developing modular SL architectures with interchangeable components tailored to various application requirements, or exploring meta learning techniques that rapidly adapt to new domains with minimum fine-tuning. By addressing these challenges, SL in EI can evolve to meet the diverse needs of future applications while maintaining its fundamental advantages in collaborative, adaptive, and secure distributed learning.
VIII-B Open Problems and Future Research
Despite significant progress in integrating SL in EI, there remain open problems in this research area, offering directions for future exploration.
VIII-B1 Strengthen Collaborative Decision-making
One of the key challenges in integrating SL with EI is the enhancement of collaborative decision-making mechanisms. Future research should aim to develop SL algorithms capable of effectively simulating complex social decision-making processes. For instance, in smart city traffic management, MARL algorithms incorporating graph neural networks could effectively model collective vehicle and pedestrian behaviors. Privacy-preserving techniques are crucial. In healthcare, for example, FL with secure aggregation protocols and differential privacy could enable collaborative model training across hospitals without compromising patient data. Adaptive meta learning approaches, such as MAML, should be investigated for rapid policy adaptation in dynamic industrial settings. These advanced decision models, which balance individual privacy with collective intelligence, could significantly enhance the effectiveness of EI systems across various domains, from optimizing energy consumption in smart grids to enhancing robotic control in manufacturing.
VIII-B2 Optimize Resource Allocation and Management
Optimizing resource allocation and management in EI using SL principles remains a critical challenge. Future research should concentrate on developing SL-inspired strategies that leverage collective intelligence to dynamically optimize resources. For instance, in large-scale IoT networks, social influence-aware FL could optimize resource allocation across heterogeneous devices. By incorporating social rules analysis techniques, this approach could model device interactions and information propagation, enabling more efficient collaborative learning and resource sharing. Another promising direction is the development of socialized MARL algorithms for adaptive edge caching. These algorithms could learn from user behavior patterns and social connections to predict content popularity and optimize cache placement, significantly reducing latency and bandwidth usage. Moreover, addressing ethical concerns in resource allocation necessitates the development of fairness-aware SL algorithms that incorporate social choice theory concepts to ensure equitable resource allocation and respect for individual device preferences and capabilities.
VIII-B3 Enhance Collaborative Learning and Knowledge Sharing
Advancing collaborative learning and information exchange within SL and EI is vital for creating an engaging and inclusive learning environment. Future research should develop advanced SL algorithms and platforms to facilitate seamless collaboration in EI. Socialized knowledge graphs incorporating social relationships and expertise levels can dynamically represent collective intelligence across edge devices, using adaptive attention mechanisms to prioritize information flow based on trust and relevance. Multi-agent transfer learning frameworks should be explored to promote active participation and diverse knowledge integration. This approach enables devices to share and adapt knowledge across domains based on social roles, using meta learning for rapid adaptation. Decentralized consensus algorithms inspired by social choice theory allow fair aggregation of diverse perspectives from edge devices while respecting individual preferences. Focusing on these SL-based algorithms and platforms can foster an engaging and inclusive learning environment, enhancing interaction and knowledge exchange within SL-enhanced EI systems.
VIII-B4 Ensuring Scalability and Interoperability
Ensuring the scalability and interoperability of SL algorithms is critical as EI systems grow in complexity and heterogeneity. Future research should develop adaptive SL frameworks that dynamically adjust to the varying system scales and capabilities of devices. For instance, it is warranted to explore hierarchical SL models where edge devices form local social clusters for efficient knowledge sharing. Higher-level nodes could manage inter-cluster communication, potentially modeled by graph neural networks. To address interoperability, universal SL interfaces leveraging semantic web technologies could facilitate seamless integration across EI platforms, using ontology-based representations of SL processes. To optimize performance in large-scale deployments, developing incremental SL algorithms that allow continuous model updates without full retraining is essential. Exploring socialized load-balancing strategies, where tasks are distributed based on social roles and connectivity patterns and dynamically adapted using the RL algorithm, is also crucial.
VIII-B5 Enhance Privacy Protection and Data Security
Addressing privacy protection and data security is crucial for integrating SL with EI. Future research should focus on devising innovative strategies to enhance privacy and security while preserving the collaborative nature of SL. These strategies should effectively safeguard personal data, employing advanced methods like context-aware encryption schemes that adapt based on the current learning phase and data importance. Exploring complex anonymization algorithms, including multi-dimensional privacy preservation techniques that protect both individual data and collective learning patterns, is necessary. Implementing robust access control mechanisms, such as adaptive multi-factor authentication systems based on the device’s learning behavior and contribution patterns, is essential. These strategies should efficiently process data in dynamic EI environments to balance utility and privacy. Ensuring EI systems implementing SL are intelligent, collaborative, secure, and privacy-preserving is crucial for building trust and encouraging wider adoption in various applications.
IX Conclusion
In this paper, we have explored the integration of SL and EI, two emerging paradigms that promise to revolutionize the fields of AI and EC. Specifically, we have discussed the background and fundamentals of SL, drawing inspiration from human social interactions and behaviors. We have also reviewed the existing concepts and paradigms related to SL, highlighting its distinctions and overlaps. Moreover, we have motivated the combination of SL and EI from two perspectives: how SL can address challenges faced by EI, and how EI can provide a suitable platform for SL implementation. To further illustrate this integration, we have presented three key aspects: socialized architecture, socialized training and socialized inference. Finally, we have investigated some future applications and tutorials of SL in EI. We have also discussed some open problems and identified future research directions on combining SL and EI. We believe that SL in EI has the potential to create more intelligent, adaptive, and robust systems that benefit various domains and applications.